A Complete Bibliography of Publications in ESAIM: Probability and Statistics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 August 2024 Version 1.15

Title word cross-reference

1 [474, 519]. 2 [31]. [0, n] [353]. α [264, 291]. β [291]. ℓ_{1} [530, 393, 313]. γ [451]. ∞ [130]. k [540, 609]. κ [280]. L^{2} [612]. L^{p} [142, 495]. L^{r} [196]. L^{s} [196]. L_{p} [242, 94, 22, 436]. M [473, 317, 243, 127]. L^{p} [541]. R^{d} [597, 301, 129]. R^{n} [144]. Z [571, 275]. Z^{+} [275]. n [129]. P [36]. p ∈ (1, 2) [436]. R^{+} [445]. σ [236]. t [151]. U [31, 102, 244, 592]. W [524]. Z [524]. Z^{2} [207].

-boosting [612]. -dependent [127]. -deviation [94]. -estimates [243]. -estimation [473, 317]. -finite [236]. -means [280]. -NN [540, 609]. -norm [530]. -optimal [196]. -oracle [393, 313].

-rate-optimality [196]. -reflected [451]. -Solutions [436]. -space [129]. -statistics [151, 592, 31, 102, 244]. -theory [242]. -time [264]. -valued [301]. -values [36].

1 [66].

2D [567].

abelian [550]. ABM [544]. abrupt [54]. absorbing [435]. absorption [214]. actuarial [148]. adaptative [569]. adapted [578]. Adaptive [75, 339, 55, 119, 462, 215, 98, 267, 315, 304, 422, 380, 40, 156, 505]. adaptivity [322]. Adding [331]. additive [383, 164, 241, 101, 324, 485, 351, 212]. Admissibility [66]. Advances [115]. affects [280]. age [179]. age-structured [179]. agents [311]. air [109]. Algebraic [525]. algorithm [227, 280, 299, 604, 558]. algorithmes [46]. Algorithms [374, 380, 288, 177, 495, 589, 285, 46]. alignment [153]. Almost [225, 288, 495, 340, 23]. along [378]. alternate [469]. alternative [331, 348]. alternatives [300]. American [61]. ANA [502]. Analysis [380, 428, 566, 174, 551, 412, 305, 309, 589, 595, 229]. Analytic [133, 20]. ancestral [486]. Anisotropic [567, 165]. Annealed [120]. anticipatifs [39]. Anticipatory [39]. Antiduality [490]. Application [426, 103, 427, 345, 299, 74, 12, 302, 489, 91, 481, 472, 440, 467, 348]. Applications [90, 27, 602, 253, 524, 170, 32, 473, 197, 292, 308, 598, 119, 101, 381, 279, 358, 286, 229, 148, 243, 8]. approach [159, 104, 326, 403, 439, 203, 584, 344, 512]. approaches [133]. Approximating [562]. Approximation [61, 26, 600, 143, 558, 539, 595, 329, 572, 390, 288, 349, 407, 79, 365, 248, 8, 555, 346, 52, 589, 560, 93, 13, 1, 341, 62, 467]. Approximations [578, 77, 27, 279, 18]. ARCH [90]. Archimedean [437]. area [515]. arithmetic [320]. arithmetically [281]. arithmétique [320]. ARMA [57]. armed [112]. arrays [206, 286]. aspects [525]. associated [526, 545, 370]. assumptions [500]. asymmetric [140]. Asymptotic [599, 413, 79, 270, 82, 96, 381, 50, 577, 59, 185, 374, 146, 337, 285, 240, 566, 388, 105, 265, 167, 2, 84, 218, 29, 553, 305, 550, 142, 612, 272, 597, 589, 282, 416, 267, 239, 575]. Asymptotically [213]. Asymptotics [456, 94, 306, 262, 561, 308, 144, 10]. atypical [109]. augmentation [234]. auto [41]. auto- [41]. Autocovariance [57]. Automata [64]. automatic [194]. autoregression [192]. autoregressive [327, 305, 447, 334, 479, 265, 335, 338]. auxiliary [523, 117]. average [474]. averaged [495]. averages [302, 23, 493]. Averaging [63, 455].

Backward [549, 123, 578, 135, 595, 436, 541]. ball [155]. balls [399, 513, 139, 267]. Banach [572, 99, 498]. bandit [112]. Bands [333, 405, 422]. bandwidth [546]. barrier [372]. barriers [135]. barycenters [574, 455]. Based [224, 322, 459, 361, 364, 152, 314, 82, 156, 347, 397, 547, 502, 565]. bases [296, 448]. basic [535]. basis [208]. Batches [66]. Bayes [178]. Bayesian [472, 227, 574, 392, 568, 336, 408, 499, 277, 441]. be [118, 341]. Beckner [482]. Behavior [149, 446, 305, 413, 577, 59, 105, 167]. Behaviour [599, 573, 605, 111, 571, 552, 79, 181, 514, 424, 179]. Bernoulli [548, 96, 275, 66]. Bernstein [511]. Berry [48, 502]. Besov [34, 80]. Bessel [389]. best [160]. beta [339]. between [23, 472, 554, 363]. bias [343]. biased [238]. Bickel [479]. biconvexity [534]. Bifractional [398]. bifurcating [395, 338]. ‘Big [5]. binary [429, 222, 433, 569]. Binomial [130, 356, 353]. Binomial-Poisson [130]. bins [118]. biological [47]. bipartite [592]. birthtimes [484]. block [407, 579, 532, 488]. blockwise [127]. bodies [34]. Boltzmann [87, 18]. boosting [612]. Bootstrapping [447, 125]. border [398]. Bound [24, 502, 467]. boundaries [95, 176]. boundary [385, 559, 510]. bounded [503, 552, 99, 166]. Bounds [29, 16, 164, 597, 555, 310, 289, 203, 363, 317, 108, 148]. Branching [222, 126, 556, 211, 494, 492, 413, 429, 484, 486, 258]. Brascamp [218]. bridge [389]. bridges [367]. Brownian [295, 520, 49, 276, 309, 143, 561, 81, 392, 162, 139, 439, 606, 294, 336, 290, 76, 106, 321, 483, 398, 278, 144, 264, 529, 370, 389, 193, 212, 444, 236]. Brownian-driven [561]. brownien [321]. Brunk [127]. BSDE [372, 424]. BSDEs [158, 349, 331, 410, 533, 603, 121]. Bulk [66].

Burgers [14].

calibration [581, 252]. call [349]. canonical [387, 519]. Capacity [203]. caravan [499]. Cardinality [547]. Carlo [469, 161, 431, 354]. CART [357]. Carthaginian [307]. case [605, 380, 139, 560, 252]. cases [398]. CDMA [203]. Cellular [64]. centered [114]. Central [365, 223, 433, 89, 170, 465, 288, 43, 272, 346, 22, 286, 518, 467]. certain [97, 90, 526, 545]. chaînes [17]. Chain [532, 6, 342, 365, 354]. Chain-referral [532]. Chains [585, 432, 383, 378, 187, 74, 485, 564, 347, 17, 395, 229, 422, 28]. change [461]. change-point [461]. changes [54]. chaos [186]. chaoses [496]. characterisation [443]. characteristic [510]. Characterization [455]. characterizing [65]. Cheeger [17]. chemical [96]. Chernoff [48]. chi [418]. chi-type [418]. Christoffel [565]. Chung [127]. circle [49, 312]. circular [21, 488]. Class [200, 90, 111, 600, 557, 249, 64, 428, 241, 539, 526, 545]. classes [544, 40, 607]. Classification [115, 131, 449]. classifier [505]. closeness [363, 543]. cloud [417]. CLT [390]. cluster [446, 96]. clustering [219, 298, 315]. clusters [152, 44]. coagulation [306]. coagulation-fragmentation [306]. coalescing [607]. codes [471]. coding [80]. coefficient [163, 600, 394, 37, 408, 14]. coefficients [123, 107, 159, 154, 397, 533, 157, 338]. Collective [374]. Collective-Rearrangement [374]. collector [419, 490]. collisions [87]. coloured [133]. comme [36]. Compact [360, 292, 560]. Comparison [116, 194, 369, 160, 593]. Competing [444, 405]. competition [371]. complete [329, 469]. Complex [477]. component [324]. components [33, 306, 122].

Composite [568]. Compound [606, 1]. compression [80]. compromise [472]. concavity [124]. Concentration [166, 383, 217]. condition [312, 564, 436, 424]. Conditional [606, 113, 475, 314, 472, 430]. Conditioned [420, 258, 341]. Conditions [147, 302]. Confidence [333, 182, 405, 267, 422]. congestion [221]. conic [12]. conjectures [120]. connected [77]. connections [264]. Consistency [376, 463, 540, 476, 350, 609, 406]. Consistent [137, 336, 459, 297, 483]. constant [548, 160, 328]. Constraints [83, 331, 410]. Construction [606, 298]. constructions [194, 13]. contact [582]. contact-mediated [582]. contaminants [384]. content [297]. context [484, 277]. continu [183]. Continuous [403, 506, 258, 556, 600, 183, 55, 294, 493]. Continuous-time [506, 258, 183, 493]. continuously [246]. continuum [583]. contraction [408, 196]. control [524, 138, 604]. controlled [365]. converge [381]. Convergence [379, 74, 100, 597, 560, 563, 322, 465, 163, 384, 288, 458, 491, 222, 11, 495, 87, 273, 92, 180, 22, 565]. convergences [399]. Convex [145, 605, 495, 294, 410, 360]. Convolution [310]. convolutions [155, 268]. coordinate [236]. coprocess [434]. copula [347]. copula-based [347]. copulas [437, 397]. Correction [545]. correlated [238]. correlations [105]. Corrigendum [158]. count [544]. counter [522]. counter-curse [522]. counting [421]. counts [407, 306, 1]. Coupling [211, 93, 584]. Couplings [576, 531]. coupon [419, 490]. covariance [379, 332, 60, 488]. covariate [375, 190]. covariates [373]. cover [557]. Cramer [564, 598, 244, 438]. Cramér-type [438]. credit [136]. criterion [239]. critical [556, 413]. Crump [433]. cup [419]. Curie [316]. curse [522]. curve [156]. curved [385, 558]. curves [543]. Cutoff [28]. Cyclic [129]. cylinder [325]. cylinders [284].

D [474, 519]. D-lattice [519]. damping [463]. Data [252, 322, 544, 41, 361, 364, 109, 156, 589, 59, 463, 18, 406, 417]. Data-driven [252, 417]. de-noising [499]. decay [440]. decaying [105]. decentralized [604]. Decomposition [566, 294]. deconvolution [178, 58]. decreasing [149]. deficiency [268]. defined [172]. Deformed [611]. degeneracy [354]. degenerate [426, 149]. degree [440]. degree-sequence [440]. delayed [372]. Densities [606, 432, 40, 310, 454, 268, 199, 192]. Density [182, 326, 278, 15, 42, 563, 339, 73, 456, 314, 55, 215, 185, 98, 267, 25, 315, 348, 150, 108, 157, 430, 283, 593]. dependence [452, 404, 110, 215, 484, 347, 397]. Dependent [170, 123, 41, 100, 142, 300, 56, 191, 536, 223, 286, 157, 328, 127, 449]. depends [298]. depth [457, 406]. derivative [353, 39]. derivatives [361, 136, 283]. dérivé [39]. descent [473]. describing [179]. design [67, 471, 85]. designed [502]. designs [361]. Detecting [54, 109]. detection [119]. determinantal [513]. Deterministic [346, 355, 497, 469, 411, 604]. Deviation [164, 498, 395, 97, 31, 494, 435, 478, 94, 4, 289, 500, 591]. Deviations [78, 21, 281, 32, 520, 151, 88, 241, 209, 101, 414, 247, 244, 203, 22, 526, 545, 103, 510, 335, 229, 316, 284, 438, 325]. d’experts [36]. Dickman [381]. difference [418]. differences [498]. different [559]. differentiability [531]. differentiable [454]. Differential [352, 123, 456, 394, 172, 431, 428, 248, 408, 259, 132, 180, 135, 105, 436, 63, 541]. diffusion [563, 385, 163, 195, 394, 94, 462, 8, 37, 52, 577, 86, 560, 246, 388, 39, 250]. diffusion-approximation [8]. Diffusions [50, 51, 72, 14, 432, 605, 402, 232, 539, 426, 181, 68, 422, 596]. Digital [186]. dilatation [196]. dimension [61, 131, 540, 346, 340, 609, 488, 401, 396]. Dimensional [576, 42, 242, 211, 402, 399, 368, 557, 561, 154, 352, 102, 416, 246, 487, 529, 68, 284, 467]. dimensionality [522]. Dimensions [24, 73]. direct [554]. Directed [220, 247]. directions [129]. Dirichlet [214, 344]. discontinuity [470]. discontinuous [352, 372, 601, 400]. discovering [47]. discovery [348]. Discrete [311, 568, 37, 148, 237, 342, 55, 91, 477, 250]. discretely [493]. discrétisation [39]. discretization [39]. disk [559]. Dislocation [255]. dispersion [544, 408]. distance [531, 597, 96]. distances [290]. Distortion [169]. distributed [351]. distribution [237, 107, 29, 70, 600, 142, 114, 443, 381, 539, 91, 560, 330, 526, 65, 341, 243, 62]. Distributions [611, 83, 390, 562, 100, 527, 575, 472, 363, 504, 545]. divergence [132, 180]. divergent [488]. divisible [128, 100, 302, 254]. division [442]. Dixie [419]. DNA [1]. does [210]. domain [52]. Dominating [226]. Donsker [154]. dormancy [582]. double [419]. Doubly [349, 436]. drift [600, 570, 468, 392, 352, 462, 426, 14, 246, 422]. Driven [181, 456, 468, 561, 248, 132, 180, 105, 595, 252, 510, 417]. dual [274]. d’une [39]. Durbin [305, 335]. Dynamical [42, 257, 74, 271, 63, 316]. dynamics [120, 365, 11, 7]. Dynamiques [120]. Dyson [49].

Edgeworth [229]. Editorial [259]. effect [227]. effects [394]. efficacy [277]. efficiencies [291]. efficiency [489, 337, 575]. Efficient [25, 471]. EGARCH [60]. eigenvalues [478, 214]. eigenvector [492]. electricity [542]. electrostatic [49]. Elementary [584]. elements [136]. ellipsoids [34]. elliptic [68]. Empirical [224, 405, 508, 59, 522, 523, 379, 142, 219, 114, 597, 173, 565]. Enhanced [197]. enlarged [359]. enlargement [307]. ensemble [519]. ensembles [387]. Entrelacements [231]. Entropic [226, 147, 130]. entropy [475, 299]. Envelope [61]. environment [391, 220, 88, 293]. environments [556]. Epidemic [586, 211]. epidemics [582]. Equal [107]. equality [425]. equation [551, 242, 73, 456, 474, 408, 527, 358, 366, 18, 10]. Equations [567, 123, 394, 172, 431, 428, 594, 248, 101, 578, 259, 14, 132, 180, 135, 105, 436, 63, 549, 541, 352]. equilibrium [491]. equipartition [270]. equivalence [559, 388]. Erased [24]. Erdos [23, 103]. Ergodic [573, 600, 539, 250]. Ergodicity [90, 426, 409, 435, 485, 486, 601]. Erlang [491]. error [238, 334, 555, 190, 595]. errors [50, 51, 502]. Esseen [502, 48]. estimate [84, 376, 431]. Estimates [224, 450, 188, 494, 95, 560, 176, 243]. Estimating [237, 542, 442, 430, 422, 20, 122, 417]. Estimation [165, 501, 334, 394, 110, 587, 190, 199, 42, 250, 269, 227, 182, 563, 391, 322, 128, 538, 361, 364, 339, 295, 104, 537, 475, 40, 375, 434, 187, 299, 534, 323, 473, 314, 55, 343, 292, 480, 326, 178, 462, 156, 215, 324, 37, 336, 408, 411, 481, 98, 267, 332, 246, 25, 315, 546, 348, 547, 317, 108, 39, 283, 117, 593, 401, 396]. estimator [238, 355, 459, 152, 356, 297, 94, 350, 463, 71, 502]. Estimators [333, 470, 51, 185, 337, 483, 216, 430, 192]. Euclidean [260]. Euler [163, 600, 52, 149, 467]. event [405, 400]. events [405]. Exact [40, 561, 352, 363]. example [385]. exceedance [251]. excesses [62]. exchange [111]. exchangeable [592]. excursion [515]. existence [312, 398]. Exit [521, 605, 6, 510, 596]. exit-problem [605]. Exit-time [521]. exit-times [596]. expanding [583]. Expansions [243, 253, 29, 262, 150, 229]. Expectation [224, 558, 452]. Expectations [38, 431]. expected [141, 597]. experimental [11, 471]. experiments [114]. expert [36]. explicit [600, 597, 389]. exploration [429]. Exponent [24]. Exponential [383, 509, 43, 173, 484, 268, 601, 507, 544, 558, 310, 514, 19]. exponents [77, 287]. Exposed [586]. exposure [384]. extension [580]. external [316]. Extinction [556, 258]. Extremal [351]. Extreme [95, 453, 478, 176]. Extremes [418, 451].

factor [271, 160]. false [348]. family [558, 514]. Farima [581]. Fast [581, 542]. fat [517]. Feller [90]. Feynman [77, 120]. field [217, 403, 584, 316, 604, 521]. fields [218, 165, 54, 260, 528, 43, 81, 515, 498, 340, 259, 261, 290, 25, 518]. Filamentous [573]. Filtering [195]. filters [272]. filtration [233, 549, 359]. Filtrations [606, 307, 234]. finies [17]. Finite [587, 572, 100, 393, 279, 519, 313, 17, 590, 199, 229, 236, 19]. finite-range [519]. finitely [550]. first [524, 305, 96, 335, 284, 328]. first-order [305, 335]. first-passage [96]. Fisher [195]. Fit [200, 300, 45, 56, 479, 575]. Fixed [291, 470, 300]. Fixed- [291]. Flow [610, 284]. flows [325, 607]. fluctuation [153]. Fold [448]. Food [585, 384]. forest [371]. forests [457, 536]. form [132, 180]. forms [218, 83, 32, 493, 344]. Formula [610, 254, 359]. formulae [570]. formulas [389]. Formulation [602, 603]. formulations [460]. Forward [123, 595]. Forward-backward [123]. fractals [249]. Fractional [276, 551, 242, 553, 537, 309, 143, 162, 464, 261, 76, 106, 168, 366, 264]. Fragmentation [573, 552, 306, 255]. Fragmentations [69]. framework [257, 277, 533]. Fréchet [312]. free [477]. freeze [396]. French [320, 231, 35, 183, 120, 321, 17, 36, 46, 39, 319]. frequency [537, 282]. Frobenius [492]. frog [550]. frontier [350]. full [18, 229]. fully [417]. function [361, 339, 501, 142, 474, 190, 117, 565]. Functional [91, 175, 361, 399, 43, 225, 485, 98, 406, 103, 344, 449]. functionals [383, 164, 342, 209, 481, 466, 25, 518, 89, 212]. functions [524, 205, 249, 405, 294]. Fungus [573]. Further [438].

G [66]. gain [470]. Galton [517, 420, 492, 395]. game [186]. games [469]. Gamma [602, 390]. gap [402, 7]. gas [87]. Gaussian [453, 29, 70, 538, 507, 32, 165, 537, 323, 260, 249, 197, 528, 451, 393, 515, 59, 340, 122, 259, 282, 168, 527, 25, 213, 315, 252, 239, 313, 160, 245, 382, 189]. Gaussianity [503]. Gelfand [231]. genealogical [371]. genealogies [378]. General [586, 341, 201, 516, 484, 289, 366, 595, 415, 549, 371, 255]. generalizations [177]. Generalized [602, 274, 443, 14, 527, 490, 358, 10, 504]. generalizing [419]. generated [249, 550, 132, 180, 351]. Generator [72]. generators [437, 372]. genomic [229]. geometric [517, 257, 91, 565]. Gibbs [81, 175]. given [232, 491]. global [322]. Goldenshluger [296]. Goodness [300, 45, 56, 200, 479, 575]. Goodness-of-Fit [200, 300, 56, 479, 575]. gradient [123, 473, 495]. gradients [91]. Gram [412]. Gramian [598]. Graph [509, 152, 298, 440]. graph-based [152]. graphical [554]. graphon [407]. graphs [44, 440]. grazing [87]. group [453, 231, 92, 569, 416]. groupes [231]. groups [550]. Growth [24, 573, 552, 484]. Growth-Fragmentation [573, 552].

half [52]. half-space [52]. Hamiltonian [463, 591]. hard [304]. hard-thresholding [304]. harmonizable [553]. Harris [289]. Hastings [299]. Hausdorff [340]. having [289]. Hawkes [491]. hazard [190]. heat [16, 242, 262]. heavy [553, 342, 308]. heavy-tailed [308]. Hedging [147, 136]. Height [371]. Heisenberg [590]. Hermite [323]. heteroscedastic [238, 448]. heuristics [448]. Hidden [391, 187, 480]. hierarchical [465, 270]. high [537, 368, 350, 282, 416, 487, 396]. high-dimensional [368, 487]. high-frequency [537, 282]. higher [192]. higher-order [192]. Hilbert [598, 213]. Hilbertian [198]. Histogram [189, 237, 118, 194, 430]. hitting [89, 167]. Hodgkin [426]. Hoeffding [235]. Hölder [287, 218, 511]. Hölderian [198]. homeomorphic [607]. homodyne [204]. homogeneous [376, 378, 232, 433, 466, 415]. Homogenization [159, 188]. homogenizations [22]. honor [230]. horizon [106]. host [582]. hull [228]. Hurst [295]. Huxley [426]. hyperbolic [260, 443, 126, 157]. hypersurfaces [518]. hypoelliptic [594]. hypotheses [75]. Hypothesis [568, 34, 348].

I.I.D. [78, 116]. idealized [380]. identifiability [330]. identification [245]. II [51, 578, 609, 180]. III [13]. image [11]. immigration [413]. Impact [457]. implied [450]. importance [497, 161]. imposed [83]. Improved [500]. incompatible [472]. incorrectly [114]. increasing [411]. increment [295]. Incremental [287]. increments [538, 451, 206]. indecomposable [413]. independent [379, 293, 330, 103]. index [364, 295, 501, 337]. indexed [249, 102, 261, 290, 13]. indices [476, 396]. inégalité [17]. inequalities [383, 450, 16, 217, 402, 130, 184, 478, 43, 387, 173, 498, 91, 416, 4, 48, 166, 395, 500, 344, 482, 175]. inequality [218, 31, 435, 393, 353, 313, 17]. Infectious [586]. Inference [107, 530, 441, 461, 172, 397, 565]. Infinite [399, 139, 444, 242, 211, 131, 96, 346, 106, 68]. infinite-dimensional [242]. infinitely [128, 100, 302, 254]. information [523, 469, 423, 140, 117]. inhomogeneous [322, 458, 184, 480, 156, 577, 336, 388]. initial [18]. initialization [280]. input [369]. inputs [561]. instrumental [377]. integral [578, 549, 236]. Integrals [253, 218, 20]. Integrated [406, 462, 37, 370]. Integration [257, 570]. Interacting [81, 605, 311, 562, 525]. interaction [217, 139, 519]. interactions [596]. interlacings [231]. Intermediate [489, 561, 575]. interpretation [87, 14]. intertwinings [477]. intervals [182, 109]. invariance [198, 511]. Invariant [525, 563, 600, 439, 539, 68, 422]. inverse [282, 527, 481, 216, 228, 304]. irregular [14]. isomorphism [440]. isotonic [45, 362]. Itô [610]. iterated [99, 278, 71, 144]. iterates [74]. Iterative [362]. IX [212].

Jagers [433]. joint [187]. jump [563, 462, 539, 411, 388, 584, 89]. jump-diffusion [563]. jumps [542, 331, 424, 549, 541].

K.P.P. [10]. Kac [77, 120]. Kantorovich [274, 511]. Kawasaki [7]. Kellerer [266]. Kernel [82, 355, 16, 262, 95, 350, 442, 71, 546, 593]. Kernel-based [82]. kernels [339, 491]. Kesten [504]. killed [341]. kind [234]. kinetic [366]. kit [524]. KMT [13]. Kolmogorov [594, 577]. Komlos [19]. KPZ [254]. Kriging [472]. Krylov [610]. Kummer [527].

Lacunar [240]. Lacunary [276]. LAMN [468]. Lancaster [506]. Langevin [346]. Laplace [555, 27, 485, 10]. Laplacians [465]. Large [21, 281, 520, 494, 101, 247, 22, 103, 510, 229, 179, 446, 97, 378, 32, 26, 31, 557, 241, 23, 486, 526, 545, 284, 127, 325, 371]. laser [119]. Lasso [313, 393, 416]. last [220]. late [258]. Latent [566]. lattice [519]. Law [557, 21, 378, 516, 23, 466, 486, 71, 103, 440, 127]. laws [128, 320, 99, 330, 370]. lead [210]. learning [523, 574, 612]. least [397]. least-squares [397]. Lemma [263, 38]. length [559, 564]. lengths [414]. Lepski [296]. letters [202]. level [292, 209, 251]. level-exceedance [251]. Lévy [599, 524, 556, 553, 572, 320, 456, 468, 145, 248, 181, 146, 282, 321, 278, 595, 15, 167, 409, 255, 319, 467]. Lévy-Driven [181, 248]. Lévy-Type [409]. Lieb [218]. Lifetime [144]. lifetimes [484]. likelihood [376, 201, 405, 480, 33, 463, 543]. Limit [571, 342, 44, 102, 72, 564, 493, 424, 251, 418, 218, 170, 465, 288, 35, 183, 43, 272, 114, 365, 346, 225, 221, 223, 433, 273, 426, 23, 176, 22, 286, 415, 518, 89, 212, 179, 467]. limites [35, 183]. limiting [370]. Lindeberg [170, 3]. lineages [486]. Linear [191, 86, 72, 83, 501, 612, 491, 76, 106, 168, 171, 216, 228, 584, 198, 304, 401]. linear-quadratic [76, 106, 168]. links [554]. Lions [531]. Lipschitz [205, 163, 600, 533, 436]. LISDLG [92]. LM [544]. Local [553, 361, 295, 233, 354, 282, 240, 212, 333, 249, 340, 264, 348, 344, 50]. localized [448]. locally [159, 12, 495]. Location [514]. log [450, 124, 254, 175]. log-concavity [124]. log-infinitely [254]. log-Sobolev [450, 175]. logarithm [99, 71]. Logarithmic [184, 16, 516, 387]. logistic [416]. lois [320]. Long [605, 60, 111, 552, 323, 300, 56, 191, 59]. long-memory [59]. long-range [300, 191]. Long-time [605, 111, 552]. Loop [24]. Loop-Erased [24]. loss [182]. Lotka [585]. Lower [24, 284]. LSI [519]. Lundberg [148]. Lundberg-type [148]. Lyapunov [77].

M [130, 578, 571, 130, 66]. M-solutions [578]. M/G/1 [66]. M/M/ [130]. Macroscopic [153]. Major [19]. management [427]. Manifold [261]. Manifolds [454, 329]. manpower [427]. many [118]. maps [455]. Marc [230]. marginal [232, 543]. market [542]. Markov [346, 383, 376, 391, 90, 355, 378, 427, 6, 164, 342, 187, 458, 435, 497, 9, 184, 74, 480, 579, 141, 365, 33, 485, 354, 202, 432, 411, 48, 347, 486, 506, 17, 584, 89, 395, 229, 604, 422, 341, 28]. Markovian [26, 311, 429]. Martingale [576, 567, 138]. martingales [35, 183, 233, 498, 166]. mass [371]. Match’ [5]. Matchings [205]. matrices [379, 49]. Matrix [566, 412, 368]. maximal [284, 325]. Maximization [558]. maximizer [274]. Maximum [480, 376, 29, 70, 206]. Maxwell [87]. McKay [504]. MCMC [93]. mean [361, 217, 312, 405, 542, 403, 358, 584, 543, 604, 521]. mean-field [584, 521]. mean-reverting [542, 358]. mean-variance [403]. meander [389]. Means [329, 280, 381]. measurable [154]. measure [222, 597, 583, 510, 422, 255, 236, 251, 175]. measure-valued [583, 251]. measurement [334, 50, 51, 190]. measurements [569]. Measures [610, 64, 308, 525, 113, 169, 273, 522, 4, 22, 85, 68, 254, 500, 482]. mechanism [411]. mechanisms [584]. media [188, 8]. mediated [582]. Meeting [293]. membrane [571]. memory [187, 323, 491, 59, 564, 60]. Metastable [181]. method [182, 295, 138, 555, 27, 569, 460, 332, 228, 63, 3, 417, 341]. methods [111, 354, 298, 369]. metric [540, 290, 609]. Metropolis [299]. Microbial [582]. Minimal [410]. Minimax [20, 434, 178, 34, 322, 303]. Minimaxity [82]. Minimizer [224]. minimizing [114]. Minimum [161]. Minkowski [297]. mismatch [169]. misspecified [250]. mixed [172, 162, 302]. Mixing [302, 55, 3]. Mixture [587, 227, 201, 412, 12, 393, 33, 252, 239, 313, 199, 108, 2]. mixtures [330, 315, 150]. Möbius [490]. Mode [84, 433, 71]. Model [41, 67, 324, 122, 204, 240, 566, 373, 391, 535, 399, 49, 550, 407, 476, 334, 110, 579, 423, 302, 221, 426, 377, 416, 98, 171, 190, 252, 239, 448, 554, 316, 430, 328, 60, 502, 179]. modeling [471]. Modelling [573]. Models [90, 509, 587, 586, 376, 201, 544, 427, 412, 570, 447, 356, 407, 12, 542, 393, 480, 172, 306, 581, 583, 313, 265, 199, 2, 532]. Moderate [151, 78, 414, 335, 316, 88, 209, 244, 203, 438, 591]. Modified [387, 66, 450]. modulated [241]. modules [235]. Moebius [482]. Moment [450, 308]. moments [446, 288, 79, 350, 287, 269, 19]. Monge [274]. monotone [425, 310]. monotonicity [111, 490]. Mons [259]. Monte [161, 469, 431, 354]. Monte-Carlo [469]. Motion [276, 295, 520, 309, 392, 162, 464, 439, 336, 411, 321, 483, 398, 278, 144, 264, 529, 370, 193]. motions [553, 129, 477, 126]. mouvement [321]. moves [469]. moving [302, 23, 493]. MRW [379]. Multi [8, 586, 573, 243, 420, 413, 271, 467]. multi-dimensional [467]. multi-factor [271]. Multi-Patch [586]. Multi-scaled [8]. Multi-Tone [243]. Multi-Type [573, 420, 413]. multiclass [505]. Multidimensional [176, 415, 19, 331, 206, 193]. Multifractal [72, 240, 254]. Multifractional [290, 295, 501, 464, 287]. multilevel [380]. multinomial [526, 545, 575]. multiple [348, 271, 440]. Multiplicative [485]. Multiscale [346]. multistable [464, 287]. multitype [258]. Multivariate [253, 345, 309, 343, 393, 292, 302, 397, 71, 593]. mutations [535]. Mycelial [573].

Nagata [540, 609]. natural [353, 514]. Nearest [131, 505]. Necessary [312]. Negative [404, 524, 356, 290]. neighbor [131, 505]. nested [471]. Network [573, 110, 203]. networked [270]. networks [592]. Neumann [512]. neural [203]. Newman [419]. NN [540, 609]. noise [551, 242, 537, 480, 133, 101, 474, 273, 181, 388, 595, 58]. noising [499]. noisy [368]. Non [90, 429, 589, 416, 518, 376, 153, 378, 163, 26, 292, 491, 597, 139, 336, 408, 122, 569, 267, 216, 239, 584, 382, 189, 502, 344]. non-adaptative [569]. Non-asymptotic [589, 416, 597, 267]. Non-binary [429, 569]. Non-central [518]. non-compact [292]. non-Gaussian [382]. non-homogeneous [376]. non-linear [491, 584]. non-Lipschitz [163]. non-local [344]. non-Markovian [26, 429]. non-parametric [336, 408]. non-randomly [502]. non-reversible [139]. non-uniqueness [153]. nonasymptotic [142]. nonlinear [428, 101, 14, 126]. Nonparametric [322, 375, 348, 283, 355, 462, 489, 34, 246, 199, 502, 449]. nonregular [114]. norm [530]. Normal [611, 253, 282, 363]. Normality [50, 240, 337, 282, 285, 2]. normalized [281, 492, 438]. note [437, 402, 509, 88, 5, 256]. Number [187, 152, 33, 381, 202]. numbers [378, 557, 23, 486, 103, 127]. Numerical [593, 46]. numériques [46].

objective [495, 472]. observation [168]. observations [250]. observed [463, 246]. obstacle [512]. occupation [222]. occurrences [202]. occurring [608]. œuvre [319]. One [537, 576, 42, 61, 402, 600, 154, 352, 102, 246, 529, 500]. One-Dimensional [42, 402, 154, 352, 246, 529]. one-sided [600, 500]. One-step [537]. onto [375]. Operator [367, 74, 481]. operators [77, 132, 180]. Optimal [563, 16, 473, 472, 531, 153, 503, 572, 455, 497, 481, 213, 85, 196, 51]. Optimality [224, 196]. Optimisation [85]. optimization [403]. option [138]. Optional [359]. Options [61]. oracle [393, 416, 313]. order [123, 305, 116, 31, 187, 350, 603, 335, 192, 269]. ordering [476]. orderings [404]. orders [588]. orientations [207]. Ornstein [143]. orthant [459]. orthant-based [459]. Orthogonal [377, 506]. outliers [530]. overfitting [593]. overlap [107]. overshoot [167].

painting [44]. paires [231]. pairs [231]. parabolic [159, 111, 512]. parameter [107, 391, 201, 538, 323, 37, 39]. parameters [470, 468, 476, 394, 110, 269]. paramétre [39]. Parametric [172, 538, 336, 408]. parametrix [460]. Pareto [351]. partial [428, 225, 168, 554, 13]. Partially [401, 481]. Particle [77, 272, 525, 560, 68, 179]. Particles [444, 49, 81, 521]. Partition [314]. Partition-based [314]. parts [570]. passage [524, 220, 96, 284, 328]. Patch [586]. patched [437]. path [217, 6, 414, 580]. paths [263, 278, 386, 245, 212]. pathspaces [81]. patterns [318]. Paul [320, 321, 319]. PCA [487]. PDE [123, 159, 264]. PDEs [512]. peacocks [345, 301]. Peccati [496]. penalisation [212]. Penalisations [193, 236]. Penalization [296, 530]. Penalized [246, 216, 458, 356, 239, 593]. penalty [252]. Penultimate [390, 62, 79]. percolation [220, 548, 96, 247, 284, 328]. perfect [509]. performance [593]. periodic [159, 119, 426, 105]. periodically [447]. periodicities [47]. Periods [586]. permutation [279]. Perron [492]. Persistence [585]. perspectives [559]. perturbation [294, 76, 106, 584]. perturbations [30, 366]. Perturbed [432, 515, 63]. pessimistic [303]. phase [64]. pick [396]. pick-freeze [396]. Piecewise [346, 355, 497, 411, 604]. piecewise-deterministic [355]. plane [360]. Plug [292, 192]. Plug-in [292, 192]. poids [35, 183]. Poincaré [402, 353, 175]. point [461, 375, 308, 594, 114, 95, 318, 176, 547, 417]. Points [226]. Pointwise [87, 40, 156]. Poisson [385, 446, 30, 375, 434, 130, 356, 407, 606, 277, 510, 421, 547, 373, 1]. pollution [109]. Pólya [65]. polymer [220]. Polynomial [289, 150, 333, 361, 323]. polynomials [377, 506]. population [394, 442, 279, 179, 371, 161]. populations [583]. portfolio [403]. positive [452]. Positivity [73, 345, 588]. Posterior [408]. potential [520, 311]. Potentials [141]. pour [35, 183, 17, 46]. Power [440, 200, 146, 150, 382]. Power-law [440]. powers [57]. précis [120]. precise [120, 27]. prediction [579, 191]. predictors [554]. Preface [134, 230, 53]. presence [582]. presenting [59]. Preservation [124]. price [137]. Prices [61]. Pricing [140, 138]. Primitive [567]. principle [97, 168, 415, 198, 511, 591]. principles [113]. prior [499, 554]. Probabilistic [570, 159, 552, 594, 14, 512, 64]. probabilités [320, 46]. probabilities [155, 507, 327, 506, 526, 545, 46, 440]. probabilitistic [87]. Probability [606, 157, 320, 79, 169, 360, 472, 269]. probability-weighted [79]. problem [605, 274, 461, 557, 187, 419, 516, 326, 178, 273, 76, 106, 168, 490, 112, 604, 243]. problems [524, 11, 489, 481, 216, 228, 584, 317, 304, 512]. procedure [93]. Process [209, 59, 72, 573, 523, 29, 70, 355, 538, 211, 305, 375, 492, 497, 456, 468, 249, 429, 114, 141, 462, 37, 92, 560, 246, 335, 389, 174, 236, 346]. Processes [599, 57, 409, 418, 453, 379, 563, 446, 524, 556, 327, 572, 20, 164, 458, 435, 323, 9, 55, 197, 501, 413, 145, 451, 300, 308, 219, 491, 241, 598, 154, 56, 302, 95, 536, 508, 433, 273, 484, 414, 146, 282, 411, 479, 287, 132, 180, 48, 289, 213, 486, 388, 105, 595, 506, 176, 351, 584, 89, 258, 15, 421, 547, 198, 13, 167, 382, 373, 283, 604, 250, 601, 341, 449, 251, 338]. product [225, 4, 68, 500]. Products [253]. progressively [359]. Projections [226]. proof [266]. propagation [8]. Properties [374, 84, 83, 270, 146, 406, 265, 60].

Property [240, 468, 310]. protection [349]. provenant [231]. proximal [177]. Proximity [38]. proxy [503]. pseudo [47]. pseudo-periodicities [47]. put [118].

quadrant [439]. quadratic

[531, 182, 218, 32, 76, 106, 168, 481, 98, 493]. Qualitative [441, 75]. quantification [477]. Quantifying [543]. Quantile [333]. quantiles [514]. Quantitative [217, 384]. quantities [493]. quantization [169, 213]. quantizers [196]. quantum [559, 204]. Quasi [435, 529, 281, 562, 560, 601]. quasi-arithmetically [281]. Quasi-ergodicity [435, 601]. Quasi-stationarity [529]. quasi-stationary [562, 560]. quasilinear [512]. quelques [46]. quenched [88]. Queue [66, 130, 561].

Rademacher [160, 496]. radii [421].

Radon [165]. Raking [523]. Raking-ratio [523]. Random [509, 249, 78, 536, 259, 171, 24, 338, 208, 153, 391, 503, 281, 67, 571, 188, 520, 116, 425, 445, 399, 562, 494, 220, 49, 54, 513, 260, 88, 232, 528, 43, 394, 428, 457, 222, 293, 8, 498, 381, 113, 306, 102, 223, 508, 414, 206, 129, 275, 23, 105, 351, 103, 286, 518, 207, 421, 547, 126, 254, 543, 440, 127, 121, 158, 400]. Randomized [396]. Randomly [585, 285, 502]. randomness [210, 318]. range [300, 56, 191, 519]. Ranked [69]. Rare [400, 6]. Rate [11, 481, 565, 556, 465, 356, 612, 408, 273, 348, 196]. rates [563, 322, 552, 495, 411, 277]. ratio [523, 201, 295, 33, 277]. RCLL [135]. Rearrangement [374, 576]. rearrangements [145]. recovery [478]. rectangle [247]. recuites [120]. recurrent [405, 508, 289]. récursifs [46]. Recursive [343, 355, 470, 46]. Reducing [596]. reduction [208, 401]. Redundancy [528]. referral [532]. refinement [438]. Reflected [372, 135, 445, 349, 451, 439, 603, 541]. reflections [331]. regime [187, 57, 33, 86, 265]. regime-switching [265]. regression [322, 41, 67, 375, 434, 296, 343, 393, 45, 109, 125, 324, 350, 362, 313, 546, 448, 189, 502]. regular [361, 118]. Regularity [548, 340, 522, 580, 500]. regularization [228]. regulator [76, 106, 168]. reinforced [562]. related [584, 493, 335]. relativistic [385]. relevant [423]. Reliability [26, 9, 110]. Reliable [579]. renewal [308]. Renormalization [92]. renormalized [529]. Rényi [23, 103]. repair [26]. repair-times [26]. Repeated [253]. Replicant [80]. representation [186, 570]. representations [492, 598, 595]. repulsion [49]. repulsive [596]. Rescaled [513]. residuals [479]. respect [476]. responses [554]. Restricted [66, 238]. restrictions [534]. result [298]. résultats [120]. Results [602, 20, 142, 120, 101, 381, 180, 19]. reversible [139]. reverting [542, 358]. revisited [49, 419, 112]. Ripley [318]. Risk [228, 317, 108, 459, 292, 332]. risks [405]. Robust [587, 530]. Robustness [441]. Rosenblatt [479, 174]. Rough [263, 580, 386]. rule [540, 609]. rules [140].

same [116]. Sample [224, 217, 151]. Sample-Based [224]. sampled [223, 493]. samples [279, 28]. sampling [161, 37, 282, 389, 532]. satisfy [564]. scale [446, 524, 538, 104, 514]. scale-space [104]. scaled [367, 8]. scaling [221]. scan [511]. scenery [508]. Schedule [66]. Schémas [39, 39]. scheme [163, 600, 428, 149]. schemes [473, 52, 213]. School [259]. Schrödinger [77, 101, 474]. sde [468, 510]. SDEs [163, 600, 595, 467]. search [186, 469, 222]. seasonal [59]. second [123, 603, 269]. Seeking [423]. segment [461, 414]. Segmentation [356, 11]. selecting [369]. selection [237, 41, 67, 368, 324, 122, 204, 98, 171, 252, 239, 448, 373, 430, 189, 357]. Self [605, 273, 281, 553, 438]. Self-interacting [605]. self-normalized [281, 438]. self-similarity [553]. Self-stabilizing [273]. Semi [538, 427, 231, 166, 9, 202]. semi-bounded [166]. Semi-group [231]. semi-groupes [231]. semi-Markov [427, 9, 202]. Semi-parametric [538]. Semigroups [184, 552, 77]. semilinear [159]. Semimartingale [294]. semimartingales [294]. seminonparametric [377]. Semiparametric [58, 190]. Sensitivities [386]. separable [598]. Separation [168]. sequence [153, 116, 202, 440]. sequences [100, 498, 47, 229, 3, 1, 127]. Sequential [568, 392, 471]. Series [240, 97, 191, 595]. Server [66]. Set [534, 13, 543]. set-indexed [13]. sets [292, 360, 547]. setting [292]. setup [348]. sewing [263]. Shannon [475]. Shao [206]. shape [107, 550, 96, 417]. Sharp [32, 368, 305]. sheet [143, 162]. Shepp [419]. short [136]. shortest [414]. shortest-path [414]. shorth [109, 125]. shot [595]. should [118]. Shrinkage [271, 233, 245]. Sided [444, 571, 600, 500]. signal [119, 245]. signals [461]. similarity [553]. simple [84, 270, 344]. simulate [154]. Simulation [248, 509, 491, 352, 52, 369, 400]. simultaneously [608]. Sinai [88]. singular [424]. singularity [104]. situation [149]. size [442]. size-structured [442]. skew [483]. skewed [390]. Slope [448]. slowly [105]. Small [446, 155, 366, 456, 306, 273, 181, 18, 15]. small-noise [273]. smooth [29, 70]. smoothed [176]. smoothers [343]. Smoothness [299, 326]. Snell [61]. Sobol [476, 337, 396]. Sobolev [450, 16, 40, 184, 387, 580, 175]. soft [560]. Sojourn [275]. solution [551, 18]. Solutions [436, 567, 352, 87, 578, 121, 158, 541]. Some [115, 136, 602, 370, 389, 378, 170, 402, 342, 570, 100, 219, 381, 489, 271, 46, 167, 607]. space [385, 551, 104, 455, 217, 99, 598, 498, 52, 129, 289, 85]. space-times [385]. spaces [572, 260, 540, 290, 80, 609, 213, 235, 511]. spacetime [210, 474]. Sparse [478, 368, 396]. spatial [73, 110, 318]. spatially [322]. SPDE [157]. SPDEs [133, 111]. Specht [235]. Spectral [412, 7, 402, 25, 228]. spectrally [524]. spectrum [379]. speeds [384]. sphere [421]. spheres [260, 482]. spider [256]. spin [68]. Splitting [607, 380, 428, 359, 400]. Spontaneous [219]. Spring [259]. Squared [531]. squares [397]. Stability [491, 27, 432, 121, 158, 111, 280, 501]. stabilizing [209, 273]. stable [155, 456, 468, 501, 464, 466]. standard [363]. standardized [206]. state [556, 221, 289]. states [187]. stationarity [529]. Stationary [260, 64, 453, 218, 29, 70, 538, 562, 55, 451, 219, 598, 86, 273, 560, 25, 590, 207, 283]. statistic [305, 295, 318, 335]. Statistical [475, 411, 47, 423, 228, 406, 117]. Statistics [487, 592, 116, 31, 151, 102, 489, 244, 526, 545, 286, 511]. Steady [221]. Stein [128, 443, 555, 527]. Step [427, 537, 149]. Stochastic [227, 262, 589, 460, 132, 180, 583, 18, 567, 123, 257, 238, 242, 288, 73, 570, 473, 456, 407, 476, 404, 394, 172, 431, 579, 588, 598, 133, 365, 138, 248, 302, 495, 408, 578, 426, 93, 285, 463, 135, 366, 22, 533, 436, 549, 179, 532, 607, 591, 541, 558, 352]. Stopping [608, 547]. strategies [296, 271]. streaming [589]. strict [503]. Strong [590, 13, 163, 350, 519, 23, 127]. strongly [495, 448, 3]. structural [201, 554]. structure [57, 347, 60, 488]. structured [442, 179]. structures [270, 306, 580]. Student [363]. Studentized [244]. studies [9, 109]. study [559, 11, 252]. sub [503]. sub-Gaussianity [503]. subfiltrations [137]. subgraph [407]. submatrix [368]. subsampling [457]. sufficient [312, 401]. sum [491, 13]. summation [124]. sums [225, 223, 415]. supercritical [548, 433]. Superexponential [599]. Superposition [72]. supersolutions [410]. supOU [302]. Supplement [230]. support [101]. suprema [141]. sure [288, 495, 225, 340, 23, 245]. surrogate [471]. Survey [115, 364, 473, 415]. Survival [327]. Swendsen [11]. Switched [585]. switching [376, 90, 57, 86, 265]. switching-regime [57]. symmetric [310, 488]. symmetry [503]. system [26, 139, 203, 521]. Systems [42, 444, 74, 525, 137, 463, 271, 63, 68, 126, 591]. Szego [218].

Tail [364, 188, 279, 459, 561, 160]. tailed [553, 308]. tails [70, 342]. Talagrand [4]. tamed [600]. tandem [561]. Taylor [262]. Tempered [464]. template [227]. temps [183]. terminal [405, 424, 121, 158]. test [201, 425, 45, 56, 33, 479, 488]. Testing [12, 568, 318, 392, 489, 34, 569, 481, 348]. Tests [82, 200, 75, 300, 119, 279, 575, 199]. their [559, 349, 177]. theorem [218, 170, 465, 288, 550, 154, 272, 346, 225, 223, 433, 266, 206, 23, 132, 580, 286, 89, 496]. Théorémes [35, 183]. theorems [418, 188, 342, 35, 183, 43, 365, 44, 102, 426, 564, 176, 22, 493, 415, 518, 212, 251]. theoretical [219, 11, 381]. Theory [115, 453, 218, 242, 292, 360]. thin [247]. three [561]. three-dimensional [561]. thresholding [304]. thresholds [171]. Tightness [507]. tilted [284, 325]. Time [232, 568, 605, 551, 111, 311, 342, 552, 183, 458, 456, 55, 548, 403, 293, 191, 536, 577, 336, 426, 372, 106, 275, 506, 264, 493, 258, 15, 167, 157, 328, 121, 158, 179, 521]. time-dependent [536]. Time-homogeneous [232]. time-inhomogeneous [458, 577, 336]. times [385, 26, 214, 590, 264, 89, 608, 596]. tomography [204]. Tone [243]. tools [47]. total [345, 588, 371]. Totally [463]. tractable [569]. traffic [221]. trajectories [601]. transfer [74, 415]. transform [165]. transformations [425]. transforms [485, 278]. Transient [207, 508]. transition [355, 64, 192]. transitions [6]. transport [274, 455]. transversal [153]. travelling [126]. tree [469, 457, 395, 255]. trees [517, 420, 186, 222, 173, 371]. trials [277]. triangular [286]. truncated [285]. Tudor [496]. Tusnady [19]. Tutte [439]. Two [24, 483, 444, 571, 425, 151, 559, 557, 73, 337, 277, 135, 112, 284]. two-dimensional [557, 284]. Two-Sided [444, 571]. Type [66, 409, 573, 418, 218, 420, 413, 120, 577, 578, 290, 244, 112, 7, 438, 430, 148, 127, 251, 496]. type-II [578].

Uhlenbeck [143]. Unbiased [431, 332, 185]. unbounded [570, 564]. undershoot [167]. Uniform [458, 350, 560, 519, 333, 572, 389]. uniformly [605]. unifying [236]. Uniqueness [68, 153, 329, 273, 175]. Unit [559]. unitary [49]. Universal [540, 609, 196]. universally [297]. unknown [481, 58]. unrestricted [7]. Upper [325, 437]. upper-patched [437]. urn [516]. Using [117, 339, 600, 109, 326, 350, 595, 448, 269, 401].

V [448]. V-Fold [448]. Vacations [66]. vaccine [277]. Validation [452]. value [453, 459, 176]. value-at-risk [459]. valued [598, 498, 301, 583, 251]. values [95, 36, 369]. vanishing [612]. Variable [546, 357, 368, 564]. Variables [78, 503, 281, 116, 425, 142, 381, 223, 377, 23, 103, 286, 369, 127, 400]. variance [503, 205, 100, 161, 94, 272, 403, 58]. Variances [38]. variant [17]. variante [17]. variate [138]. variation [572, 382]. variations [146]. variogram [538]. vector [122]. vectorial [35, 183]. vectorielles [35, 183]. vectors [351, 19]. version [93, 19]. versus [296]. Very [517, 258]. VI [193]. via [237, 435, 143, 263, 161, 555, 87, 580, 386, 13, 543]. vibrometry [119]. view [552, 594]. virus [582]. Viscosity [567]. VLMC [173]. volatility [570, 468, 138, 302]. Volterra [585, 578, 549]. votes [36].

Wald [487]. Walk

[24, 391, 445, 88, 102, 206, 275, 207]. walks [571, 562, 494, 222, 256, 293, 508, 590]. Wang [11]. warped [296]. Wasserstein [531, 574, 455, 597]. Watson [517, 305, 420, 492, 335, 395]. wave [551, 73, 8, 474]. Wavelet [551, 323, 309, 240, 326, 499, 502]. wavelets [104]. waves [126]. Weak [378, 215, 581]. Weakly [598, 100, 286, 449]. Weibull [107]. Weighted [397, 402, 513, 79, 381, 113, 489, 511, 269, 35, 183]. Weiss [316]. Wellposedness [603]. Where [210]. which [381, 564]. white [551, 474, 388]. Wiener [367, 236]. without [201]. Wong [132]. word [202, 1]. work [319]. Wright [195].

Yor [230]. Young [218, 22].

Zakai [132]. zero [122, 382]. Zipf [516].

References

Schbath:1997:CPA

[1] Sophie Schbath. Compound Poisson approximation of word counts in DNA sequences. ESAIM: Probability and Statistics, 1(??):1–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

VanDeGeer:1997:ANM

[2] Sara Van De Geer. Asymptotic normality in mixture models. ESAIM: Probability and Statistics, 1(??):17–??, ????

1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rio:1997:ALM

[3] Emmanuel Rio. About the Lindeberg method for strongly mixing sequences. ESAIM: Probability and Statistics, 1 (??):35–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ledoux:1997:TDI

[4] Michel Ledoux. On Talagrand’s deviation inequalities for product measures. ESAIM: Probability and Statistics, 1 (??):63–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coulomb:1997:NBM

[5] Jean-Michel Coulomb. A note on ‘Big Match’. ESAIM: Probability and Statistics, 1(??):89–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Catoni:1997:EPM

[6] Olivier Catoni and Raphaël Cerf. The exit path of a Markov chain with rare transitions. ESAIM: Probability and Statistics, 1(??):95–??, ???? 1997. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Posta:1997:SGU

[7] Gustavo Posta. Spectral gap for an unrestricted Kawasaki type dynamics. ESAIM: Probability and Statistics, 1 (??):145–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Garnier:1997:MSD

[8] Josselin Garnier. Multi-scaled diffusionapproximation. Applications to wave propagation in random media. ESAIM: Probability and Statistics, 1(??):183–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cocozza-Thivent:1997:SMP

[9] Christiane Cocozza-Thivent and Michel Roussignol. Semi-Markov processes for reliability studies. ESAIM: Probability and Statistics, 1(??):207–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rouques:1997:LAG

[10] Jean-Philippe Rouqués. Laplace asymptotics for generalized K.P.P. equation. ESAIM: Probability and Statistics, 1 (??):225–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gaudron:1997:RCS

[11] Isabelle Gaudron. Rate of convergence of the Swendsen–Wang dynamics in image segmentation problems: a theoretical and experimental study. ESAIM: Probability and Statistics, 1(??):259–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dacunha-Castelle:1997:TLC

[12] Didier Dacunha-Castelle and Elisabeth Gassiat. Testing in locally conic models, and application to mixture models. ESAIM: Probability and Statistics, 1(??):285–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rio:1997:SAS

[13] E. Rio. Strong approximation for setindexed partial sum processes via KMT constructions III. ESAIM: Probability and Statistics, 1(??):319–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Jourdain:1997:DNI

[14] Benjamin Jourdain. Diffusions with a nonlinear irregular drift coefficient and probabilistic interpretation of generalized Burgers’ equations. ESAIM: Probability and Statistics, 1(??):339–??, ???? 1997. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Picard:1997:DST

[15] Jean Picard. Density in small time for Lévy processes. ESAIM: Probability and Statistics, 1(??):357–??, ???? 1997. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bakry:1997:OHK

[16] Dominique Bakry, Daniel Concordet, and Michel Ledoux. Optimal heat kernel bounds under logarithmic Sobolev inequalities. ESAIM: Probability and Statistics, 1(??):391–??, ???? 1997. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Miclo:1998:VIC

[17] Laurent Miclo. Une variante de l’inégalité de Cheeger pour les chaînes de Markov finies. (French) [A variant of Cheeger’s inequality for finite Markov chains]. ESAIM: Probability and Statistics, 2(??):1–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Meleard:1998:SAS

[18] Sylvie Meleard. Stochastic approximations of the solution of a full Boltzmann equation with small initial data. ESAIM: Probability and Statistics, 2 (??):23–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Zaitsev:1998:MVR

[19] Andrei Yu. Zaitsev. Multidimensional version of the results of Komlos, Major and Tusnady for vectors with finite exponential moments. ESAIM: Probability and Statistics, 2(??):41–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Benhenni:1998:MRE

[20] Karim Benhenni and Jacques Istas. Minimax results for estimating integrals of analytic processes. ESAIM: Probability and Statistics, 2(??):109–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Arous:1998:LDC

[21] Gérard Ben Arous and Ofer Zeitouni. Large deviations from the circular law. ESAIM: Probability and Statistics, 2 (??):123–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Michel:1998:LDC

[22] Julien Michel and Didier Piau. Large deviations, central limit theorems and L_{p} convergence for Young measures and stochastic homogenizations. ESAIM: Probability and Statistics, 2(??):135–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lanzinger:1998:ASL

[23] Hartmut Lanzinger. An almost sure limit theorem for moving averages of random variables between the strong law of large numbers and the Erdős–Rényi law. ESAIM: Probability and Statistics, 2(??):163–??, ???? 1998. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lawler:1999:LBG

[24] Gregory F. Lawler. A lower bound on the growth exponent for loop-erased random walk in two dimensions. ESAIM: Probability and Statistics, 3(??):1–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ludena:1999:EEF

[25] Carenne Ludeña. Efficient estimation of functionals of the spectral density of stationary Gaussian fields. ESAIM: Probability and Statistics, 3(??):23–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bon:1999:ARL

[26] Jean-Louis Bon and Jean Bretagnolle. Approximation of reliability for a large system with non-Markovian repairtimes. ESAIM: Probability and Statistics, 3(??):49–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guionnet:1999:SPL

[27] A. Guionnet. Stability of precise Laplace’s method under approximations; applications. ESAIM: Probability and Statistics, 3(??):67–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ycart:1999:CSM

[28] Bernard Ycart. Cutoff for samples of Markov chains. ESAIM: Probability and Statistics, 3(??):89–??, ???? 1999. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Azais:1999:BAE

[29] Jean-Marc Azaïs, Christine Cierco-Ayrolles, and Alain Croquette. Bounds and asymptotic expansions for the distribution of the maximum of a smooth stationary Gaussian process. ESAIM: Probability and Statistics, 3(??):107–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Barbour:1999:PP

[30] Andrew D. Barbour and Aihua Xia. Poisson perturbations. ESAIM: Probability and Statistics, 3(??):131–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bretagnolle:1999:NLD

[31] Jean Bretagnolle. A new large deviation inequality for U-statistics of order 2. ESAIM: Probability and Statistics, 3 (??):151–??, ???? 1999. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bercu:2000:SLD

[32] Bernard Bercu, Fabrice Gamboa, and Marc Lavielle. Sharp large deviations for Gaussian quadratic forms with applications. ESAIM: Probability and Statistics, 4(??):1–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gassiat:2000:LRT

[33] Elisabeth Gassiat and Christine Keribin. The likelihood ratio test for the number of components in a mixture with Markov regime. ESAIM: Probability and Statistics, 4(??):25–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ingster:2000:MNH

[34] Yuri I. Ingster and Irina A. Suslina. Minimax nonparametric hypothesis testing for ellipsoids and Besov bodies. ESAIM: Probability and Statistics, 4(??):53–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chaabane:2000:TLA

[35] Faouzi Chaabane and Faïza Maaouia. Théorémes limites avec poids pour les martingales vectorielles. (French) [Weighted limit theorems for vectorial martingales]. ESAIM: Probability and Statistics, 4(??):137–??, ???? 2000. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Morel:2000:VCV

[36] Guy Morel. Les P -values comme votes d’experts. (French) [P -values as expert votes]. ESAIM: Probability and Statistics, 4(??):191–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gloter:2000:DSI

[37] Arnaud Gloter. Discrete sampling of an integrated diffusion process and parameter estimation of the diffusion coefficient. ESAIM: Probability and Statistics, 4(??):205–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sonsino:2000:LPV

[38] Doron Sonsino. A lemma on proximity of variances and expectations. ESAIM: Probability and Statistics, 4(??):229–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Samos:2000:SDA

[39] Sandie Souchet Samos. Schémas de discrétisation anticipatifs et estimation du paramétre de dérivé d’une diffusion. (French) [Anticipatory discretization schemas en parameter estimation of the derivative of a diffusion]. ESAIM: Probability and Statistics, 4(??):233–??, ???? 2000. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Butucea:2001:EAP

[40] Cristina Butucea. Exact adaptive pointwise estimation on Sobolev classes of densities. ESAIM: Probability and Statistics, 5(??):1–??, ???? 2001. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baraud:2001:MSA

[41] Yannick Baraud, F. Comte, and G. Viennet. Model selection for (auto-)regression with dependent data. ESAIM: Probability and Statistics, 5 (??):33–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Prieur:2001:DEO

[42] Clémentine Prieur. Density estimation for one-dimensional dynamical systems. ESAIM: Probability and Statistics, 5 (??):51–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dedecker:2001:EIF

[43] Jérôme Dedecker. Exponential inequalities and functional central limit theorems for random fields. ESAIM: Probability and Statistics, 5(??):77–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Garet:2001:LTP

[44] Olivier Garet. Limit theorems for the painting of graphs by clusters. ESAIM: Probability and Statistics, 5(??):105–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Durot:2001:GFT

[45] Cécile Durot and Anne-Sophie Tocquet. Goodness of fit test for isotonic regression. ESAIM: Probability and Statistics, 5(??):119–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pages:2001:QAR

[46] Gilles Pagés. Sur quelques algorithmes récursifs pour les probabilités numériques. (French) [On some recursive algorithms for numerical probabilities]. ESAIM: Probability and Statistics, 5(??):141–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Prum:2001:STD

[47] Bernard Prum, Élisabeth de Turckheim, and Martin Vingron. Statistical tools for discovering pseudo-periodicities in biological sequences. ESAIM: Probability and Statistics, 5(??):171–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lezaud:2001:CBE

[48] Pascal Lezaud. Chernoff and Berry– Esséen inequalities for Markov processes. ESAIM: Probability and Statistics, 5(??):183–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cepa:2001:BPE

[49] Emmanuel Cépa and Dominique Lépingle. Brownian particles with electrostatic repulsion on the circle: Dyson’s model for unitary random matrices revisited. ESAIM: Probability and Statistics, 5 (??):203–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gloter:2001:DMEa

[50] Arnaud Gloter and Jean Jacod. Diffusions with measurement errors. I. Local asymptotic normality. ESAIM: Probability and Statistics, 5(??):225–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gloter:2001:DMEb

[51] Arnaud Gloter and Jean Jacod. Diffusions with measurement errors. II. Optimal estimators. ESAIM: Probability and Statistics, 5(??):243–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gobet:2001:ESH

[52] Emmanuel Gobet. Euler schemes and half-space approximation for the simulation of diffusion in a domain. ESAIM: Probability and Statistics, 5(??):261–??, ???? 2001. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Soulier:2002:P

[53] Phillipe Soulier. Preface. ESAIM: Probability and Statistics, 6(??):185–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chambaz:2002:DAC

[54] Antoine Chambaz. Detecting abrupt changes in random fields. ESAIM: Probability and Statistics, 6(??):189–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Comte:2002:AES

[55] Fabienne Comte and Florence Merlevéde. Adaptive estimation of the stationary density of discrete and continuous time mixing processes. ESAIM: Probability and Statistics, 6(??):211–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fay:2002:GFT

[56] Gilles Fay and Anne Philippe. Goodnessof-fit test for long range dependent processes. ESAIM: Probability and Statistics, 6(??):239–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Francq:2002:ASP

[57] Christian Francq and Jean-Michel Zakoïan. Autocovariance structure of powers of switching-regime ARMA processes. ESAIM: Probability and Statistics, 6(??):259–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Matias:2002:SDU

[58] Catherine Matias. Semiparametric deconvolution with unknown noise variance. ESAIM: Probability and Statistics, 6(??):271–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Haye:2002:ABE

[59] Mohamedou Ould Haye. Asymptotic behavior of the empirical process for Gaussian data presenting seasonal longmemory. ESAIM: Probability and Statistics, 6(??):293–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Surgailis:2002:LMP

[60] Donatas Surgailis and Marie-Claude Viano. Long memory properties and covariance structure of the EGARCH model. ESAIM: Probability and Statistics, 6(??):311–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bally:2002:ASE

[61] Vlad Bally and Bruno Saussereau. Approximation of the Snell Envelope and American Options Prices in dimension one. ESAIM: Probability and Statistics, 6(??):1–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Worms:2002:PAD

[62] Rym Worms. Penultimate approximation for the distribution of the excesses. ESAIM: Probability and Statistics, 6 (??):21–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pene:2002:AMD

[63] Françoise Péne. Averaging method for differential equations perturbed by dynamical systems. ESAIM: Probability and Statistics, 6(??):33–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

DaiPra:2002:SMP

[64] Paolo Dai Pra, Pierre-Yves Louis, and Sylvie Rœlly. Stationary measures and phase transition for a class of Probabilistic Cellular Automata. ESAIM: Probability and Statistics, 6(??):89–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ramos:2002:CPD

[65] Héctor M. Ramos, David Almorza, and Juan A. García-Ramos. On characterizing the pólya distribution. ESAIM: Probability and Statistics, 6(??):105–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Madan:2002:RAB

[66] Kailash C. Madan and Walid Abu-Dayyeh. Restricted admissibility of batches into an M/G/1 type bulk queue with modified Bernoulli schedule server vacations. ESAIM: Probability and Statistics, 6(??):113–??, ???? 2002. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baraud:2002:MSR

[67] Yannick Baraud. Model selection for regression on a random design. ESAIM: Probability and Statistics, 6(??):127–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ramirez:2002:UIP

[68] Alejandro F. Ramírez. Uniqueness of invariant product measures for elliptic infinite dimensional diffusions and particle spin systems. ESAIM: Probability and Statistics, 6(??):147–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Berestycki:2002:RF

[69] Julien Berestycki. Ranked fragmentations. ESAIM: Probability and Statistics, 6(??):157–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Azais:2002:TDM

[70] Jean-Marc Azaïs, Jean-Marc Bardet, and Mario Wschebor. On the tails of the distribution of the maximum of a smooth stationary Gaussian process. ESAIM: Probability and Statistics, 6 (??):177–??, ???? 2002. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mokkadem:2003:LIL

[71] Abdelkader Mokkadem and Mariane Pelletier. The law of the iterated logarithm for the multivariate kernel mode estimator. ESAIM: Probability and Statistics, 7(??):1–??, March 2003. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Igloi:2003:SDL

[72] Endre Iglói and György Terdik. Superposition of diffusions with linear generator and its multifractal limit process. ESAIM: Probability and Statistics, 7(??):23–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chaleyat–Maurel:2003:PDS

[73] Mireille Chaleyat-Maurel and Marta Sanz-Solé. Positivity of the density for the stochastic wave equation in two spatial dimensions. ESAIM: Probability and Statistics, 7(??):89–??, March 2003. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Conze:2003:CIT

[74] Jean-Pierre Conze and Albert Raugi. Convergence of iterates of a transfer operator, application to dynamical systems and to Markov chains. ESAIM: Probability and Statistics, 7(??):115–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baraud:2003:ATQ

[75] Yannick Baraud, Sylvie Huet, and Béatrice Laurent. Adaptive tests of qualitative hypotheses. ESAIM: Probability and Statistics, 7(??):147–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kleptsyna:2003:ALQ

[76] M. L. Kleptsyna, Alain Le Breton, and M. Viot. About the linear-quadratic regulator problem under a fractional Brownian perturbation. ESAIM: Probability and Statistics, 7(??):161–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

DelMoral:2003:PAL

[77] Pierre Del Moral and L. Miclo. Particle approximations of Lyapunov exponents connected to Schrödinger operators and Feynman–Kac semigroups. ESAIM: Probability and Statistics, 7 (??):171–??, March 2003. CODEN ????

ISSN 1292-8100 (print), 1262-3318 (electronic).

Eichelsbacher:2003:MDD

[78] Peter Eichelsbacher and Matthias Löwe. Moderate deviations for I.I.D. random variables. ESAIM: Probability and Statistics, 7(??):209–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Diebolt:2003:ABP

[79] Jean Diebolt, Armelle Guillou, and Rym Worms. Asymptotic behaviour of the probability-weighted moments and penultimate approximation. ESAIM: Probability and Statistics, 7(??):219–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kerkyacharian:2003:RCC

[80] Gérard Kerkyacharian and Dominique Picard. Replicant compression coding in Besov spaces. ESAIM: Probability and Statistics, 7(??):239–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dereudre:2003:IBP

[81] David Dereudre. Interacting Brownian particles and Gibbs fields on pathspaces. ESAIM: Probability and Statistics, 7 (??):251–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ermakov:2003:AMK

[82] Michael Ermakov. On asymptotic minimaxity of kernel-based tests. ESAIM: Probability and Statistics, 7(??):279–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Belomestny:2003:CDI

[83] Denis Belomestny. Constraints on distributions imposed by properties of linear forms. ESAIM: Probability and Statistics, 7(??):313–??, March 2003. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Abraham:2004:APS

[84] Christophe Abraham, Gérard Biau, and Benoît Cadre. On the asymptotic properties of a simple estimate of the mode. ESAIM: Probability and Statistics, 8 (??):1–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Molchanov:2004:OSM

[85] Ilya Molchanov and Sergei Zuyev. Optimisation in space of measures and optimal design. ESAIM: Probability and Statistics, 8(??):12–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guyon:2004:LDS

[86] Xavier Guyon, Serge Iovleff, and Jian-Feng Yao. Linear diffusion with stationary switching regime. ESAIM: Probability and Statistics, 8(??):25–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guerin:2004:PCB

[87] Héléne Guérin. Pointwise convergence of Boltzmann solutions for grazing collisions in a Maxwell gas via a probabilitistic interpretation. ESAIM: Probability and Statistics, 8(??):36–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Comets:2004:NQM

[88] Francis Comets and Serguei Popov. A note on quenched moderate deviations for Sinai’s random walk in random environment. ESAIM: Probability and Statistics, 8(??):56–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Paroissin:2004:CLT

[89] Christian Paroissin and Bernard Ycart. Central limit theorem for hitting times of functionals of Markov jump processes. ESAIM: Probability and Statistics, 8 (??):66–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Attali:2004:ECC

[90] Jean-Gabriel Attali. Ergodicity of a certain class of non Feller models: Applications to ARCH and Markov switching models. ESAIM: Probability and Statistics, 8(??):76–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Joulin:2004:FID

[91] Aldéric Joulin and Nicolas Privault. Functional inequalities for discrete gradients and application to the geometric distribution. ESAIM: Probability and Statistics, 8(??):87–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Igloi:2004:RGC

[92] Endre Iglói. Renormalization group of and convergence to the LISDLG process. ESAIM: Probability and Statistics, 8(??):102–??, August 2004. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kuhn:2004:CSA

[93] Estelle Kuhn and Marc Lavielle. Coupling a stochastic approximation version of EM with an MCMC procedure. ESAIM: Probability and Statistics, 8 (??):115–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Doukhan:2004:ADV

[94] Paul Doukhan and José R. León. Asymptotics for the L_{p}-deviation of the variance estimator under diffusion. ESAIM: Probability and Statistics, 8 (??):132–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Girard:2004:EVK

[95] Stéphane Girard and Pierre Jacob. Extreme values and kernel estimates of point processes boundaries. ESAIM: Probability and Statistics, 8(??):150–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Garet:2004:ASC

[96] Olivier Garet and Régine Marchand. Asymptotic shape for the chemical distance and first-passage percolation on the infinite Bernoulli cluster. ESAIM: Probability and Statistics, 8(??):169–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Arcones:2004:LDP

[97] Miguel A. Arcones. The large deviation principle for certain series. ESAIM: Probability and Statistics, 8(??):200–??, August 2004. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Laurent:2005:AEQ

[98] Béatrice Laurent. Adaptive estimation of a quadratic functional of a density by model selection. ESAIM: Probability and Statistics, 9(??):1–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Deng:2005:BLI

[99] Dianliang Deng. On the bounded laws of iterated logarithm in Banach space. ESAIM: Probability and Statistics, 9 (??):19–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dedecker:2005:CID

[100] Jérôme Dedecker and Sana Louhichi. Convergence to infinitely divisible distributions with finite variance for some weakly dependent sequences. ESAIM: Probability and Statistics, 9(??):38–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gautier:2005:LDS

[101] Éric Gautier. Large deviations and support results for nonlinear Schrödinger equations with additive noise and applications. ESAIM: Probability and Statistics, 9(??):74–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guillotin-Plantard:2005:LTS

[102] Nadine Guillotin-Plantard and Véronique Ladret. Limit theorems for U-statistics indexed by a one dimensional random walk. ESAIM: Probability and Statistics, 9(??):98–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Najim:2005:LDI

[103] Jamal Najim. Large deviations for independent random variables – application to Erdős–Rényi’s functional law of large numbers. ESAIM: Probability and Statistics, 9(??):116–??, June 2005. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bigot:2005:SSA

[104] Jérémie Bigot. A scale-space approach with wavelets to singularity estimation. ESAIM: Probability and Statistics, 9 (??):143–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marty:2005:ABD

[105] Renaud Marty. Asymptotic behavior of differential equations driven by periodic and random processes with slowly decaying correlations. ESAIM: Probability and Statistics, 9(??):165–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kleptsyna:2005:ITH

[106] Marina L. Kleptsyna, Alain Le Breton, and Michel Viot. On the infinite time horizon linear-quadratic regulator problem under a fractional Brownian perturbation. ESAIM: Probability and Statistics, 9(??):185–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Al-Saidy:2005:IOC

[107] Obaid Al-Saidy, Hani M. Samawi, and Mohammad F. Al-Saleh. Inference on overlap coefficients under the Weibull distribution: Equal shape parameter. ESAIM: Probability and Statistics, 9 (??):206–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rakhlin:2005:RBM

[108] Alexander Rakhlin, Dmitry Panchenko, and Sayan Mukherjee. Risk bounds for mixture density estimation. ESAIM: Probability and Statistics, 9(??):220–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Durot:2005:DAD

[109] Cécile Durot and Karelle Thiébot. Detecting atypical data in air pollution studies by using shorth intervals for regression. ESAIM: Probability and Statistics, 9(??):230–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dinwoodie:2005:EPN

[110] Ian Hepburn Dinwoodie. Estimation of parameters in a network reliability model with spatial dependence. ESAIM: Probability and Statistics, 9(??):241–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Berge:2005:LTB

[111] Benjamin Bergé and Bruno Saussereau. On the long-time behaviour of a class of parabolic SPDE’s: monotonicity methods and exchange of stability. ESAIM: Probability and Statistics, 9(??):254–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pages:2005:TAB

[112] Gilles Pagés. A two armed bandit type problem revisited. ESAIM: Probability and Statistics, 9(??):277–??, June 2005.

CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gozlan:2005:CPR

[113] Nathael Gozlan. Conditional principles for random weighted measures. ESAIM: Probability and Statistics, 9(??):283–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ferger:2005:MPI

[114] Dietmar Ferger. On the minimizing point of the incorrectly centered empirical process and its limit distribution in nonregular experiments. ESAIM: Probability and Statistics, 9(??):307–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boucheron:2005:TCS

[115] Stéphane Boucheron, Olivier Bousquet, and Gábor Lugosi. Theory of classification: a survey of some recent advances. ESAIM: Probability and Statistics, 9(??):323–??, June 2005. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bon:2006:COS

[116] Jean-Louis Bon and Eugen Păltănea.

Comparison of order statistics in a random sequence to the same statistics with i.i.d. variables. ESAIM: Probability and Statistics, 10(??):1–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tarima:2006:UAI

[117] Sergey Tarima and Dmitri Pavlov. Using auxiliary information in statistical function estimation. ESAIM: Probability and Statistics, 10(??):11–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Birge:2006:HMB

[118] Lucien Birgé and Yves Rozenholc. How many bins should be put in a regular histogram. ESAIM: Probability and Statistics, 10(??):24–??, September 2006. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fromont:2006:ATP

[119] Magalie Fromont and Céline Lévy-leduc. Adaptive tests for periodic signal detection with applications to laser vibrometry. ESAIM: Probability and Statistics, 10(??):46–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

DelMoral:2006:DRT

[120] Pierre Del Moral and Laurent Miclo. Dynamiques recuites de type Feynman–Kac: résultats précis et conjectures. (French) [Annealed dynamics of Feynman–Kac type: precise results and conjectures]. ESAIM: Probability and Statistics, 10(??):76–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Toldo:2006:SSB

[121] Sandrine Toldo. Stability of solutions of BSDEs with random terminal time. ESAIM: Probability and Statistics, 10 (??):141–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic). See corrigendum [158].

Huet:2006:MSE

[122] Sylvie Huet. Model selection for estimating the non zero components of a Gaussian vector. ESAIM: Probability and Statistics, 10(??):164–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Abraham:2006:FBS

[123] Romain Abraham and Olivier Riviere.

Forward-backward stochastic differential equations and PDE with gradient dependent second order coefficients. ESAIM: Probability and Statistics, 10 (??):184–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Johnson:2006:PLC

[124] Oliver Johnson and Christina Goldschmidt. Preservation of log-concavity on summation. ESAIM: Probability and Statistics, 10(??):206–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Durot:2006:BSR

[125] Cécile Durot and Karelle Thiébot. Bootstrapping the shorth for regression. ESAIM: Probability and Statistics, 10 (??):216–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ratanov:2006:BRM

[126] Nikita Ratanov. Branching random motions, nonlinear hyperbolic systems and travelling waves. ESAIM: Probability and Statistics, 10(??):236–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Thanh:2006:BCT

[127] Le Van Thanh. On the Brunk–Chung type strong law of large numbers for sequences of blockwise m-dependent random variables. ESAIM: Probability and Statistics, 10(??):258–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Averkamp:2006:SEI

[128] R. Averkamp and C. Houdré. Stein estimation for infinitely divisible laws. ESAIM: Probability and Statistics, 10 (??):269–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lachal:2006:CRM

[129] Aimé Lachal. Cyclic random motions in R^{d}-space with n directions. ESAIM: Probability and Statistics, 10(??):277– ??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chafai:2006:BPE

[130] Djalil Chafaï. Binomial-Poisson entropic inequalities and the M/M/∞ queue. ESAIM: Probability and Statistics, 10 (??):317–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cerou:2006:NNC

[131] Frédéric Cérou and Arnaud Guyader. Nearest neighbor classification in infinite dimension. ESAIM: Probability and Statistics, 10(??):340–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lejay:2006:SDE

[132] Antoine Lejay. Stochastic differential equations driven by processes generated by divergence form operators I: a Wong– Zakai theorem. ESAIM: Probability and Statistics, 10(??):356–??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ferrante:2006:SCN

[133] Marco Ferrante and Marta Sanz-Solé.

SPDEs with coloured noise: Analytic and stochastic approaches. ESAIM: Probability and Statistics, 10(??):380– ??, September 2006. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cont:2007:P

[134] Rama Cont, Jean-Pierre Fouque, and Bernard Lapeyre. Preface. ESAIM: Probability and Statistics, 11(??):1– ??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lepeltier:2007:RBS

[135] Jean-Pierre Lepeltier and Mingyu Xu.

Reflected backward stochastic differential equations with two RCLL barriers. ESAIM: Probability and Statistics, 11 (??):3–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Durand:2007:SSE

[136] Philippe Durand and Jean-Frédéric Jouanin. Some short elements on hedging credit derivatives. ESAIM: Probability and Statistics, 11(??):23–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gombani:2007:CPS

[137] Andrea Gombani, Stefan Jaschke, and Wolfgang Runggaldier. Consistent price systems for subfiltrations. ESAIM: Probability and Statistics, 11(??):35– ??, February 2007. CODEN ????

ISSN 1292-8100 (print), 1262-3318 (electronic).

Fouque:2007:MCV

[138] Jean-Pierre Fouque and Chuan-Hsiang Han. A martingale control variate method for option pricing with stochastic volatility. ESAIM: Probability and Statistics, 11(??):40–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fradon:2007:ISB

[139] Myriam Fradon and Sylvie Rœlly. Infinite system of Brownian balls with interaction: the non-reversible case. ESAIM: Probability and Statistics, 11 (??):55–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ogawa:2007:PRU

[140] Shigeyoshi Ogawa and Monique Pontier. Pricing rules under asymmetric information. ESAIM: Probability and Statistics, 11(??):80–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Follmer:2007:PMP

[141] Hans Föllmer and Thomas Knispel. Potentials of a Markov process are expected suprema. ESAIM: Probability and Statistics, 11(??):89–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dedecker:2007:EDF

[142] Jérôme Dedecker and Florence Merlevéde. The empirical distribution function for dependent variables: asymptotic and nonasymptotic results in L^{p}. ESAIM: Probability and Statistics, 11 (??):102–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coutin:2007:AFB

[143] Laure Coutin and Monique Pontier. Approximation of the fractional Brownian sheet via Ornstein–Uhlenbeck sheet. ESAIM: Probability and Statistics, 11 (??):115–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nane:2007:LAI

[144] Erkan Nane. Lifetime asymptotics of iterated Brownian motion in R^{n}. ESAIM: Probability and Statistics, 11(??):147– ??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Davydov:2007:CRL

[145] Youri Davydov and Emmanuel Thilly.

Convex rearrangements of Lévy processes. ESAIM: Probability and Statistics, 11(??):161–??, February 2007. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Jacod:2007:APP

[146] Jean Jacod. Asymptotic properties of power variations of Lévy processes. ESAIM: Probability and Statistics, 11 (??):173–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Njoh:2007:ECH

[147] Samuel Njoh. Entropic conditions and hedging. ESAIM: Probability and Statistics, 11(??):197–??, February 2007. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sendova:2007:DLT

[148] Kristina Sendova. Discrete Lundbergtype bounds with actuarial applications. ESAIM: Probability and Statistics, 11 (??):217–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lemaire:2007:BES

[149] Vincent Lemaire. Behavior of the Euler scheme with decreasing step in a degenerate situation. ESAIM: Probability and Statistics, 11(??):236–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pommeret:2007:PED

[150] Denys Pommeret. Polynomial expansions of density of power mixtures. ESAIM: Probability and Statistics, 11 (??):248–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cao:2007:MDT

[151] Hongyuan Cao. Moderate deviations for two sample t-statistics. ESAIM: Probability and Statistics, 11(??):264– ??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Biau:2007:GBE

[152] Gérard Biau, Benoît Cadre, and Bruno Pelletier. A graph-based estimator of the number of clusters. ESAIM: Probability and Statistics, 11(??):272–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Amsalu:2007:MNU

[153] Saba Amsalu, Heinrich Matzinger, and Serguei Popov. Macroscopic nonuniqueness and transversal fluctuation in optimal random sequence alignment. ESAIM: Probability and Statistics, 11 (??):281–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Etore:2007:DTS

[154] Pierre Étoré and Antoine Lejay. A Donsker theorem to simulate onedimensional processes with measurable coefficients. ESAIM: Probability and Statistics, 11(??):301–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Aurzada:2007:SBP

[155] Frank Aurzada and Thomas Simon.

Small ball probabilities for stable convolutions. ESAIM: Probability and Statistics, 11(??):327–??, February 2007. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gaiffas:2007:PAC

[156] Stéphane Gaïffas. On pointwise adaptive curve estimation based on inhomogeneous data. ESAIM: Probability and Statistics, 11(??):344–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sanz-Sole:2007:PDH

[157] Marta Sanz-Solé and Iván Torrecilla-Tarantino. Probability density for a hyperbolic SPDE with time dependent coefficients. ESAIM: Probability and Statistics, 11(??):365–??, February 2007.

CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Toldo:2007:CSS

[158] Sandrine Toldo. Corrigendum to “Stability of solutions of BSDEs with random terminal time”. ESAIM: Probability and Statistics, 11(??):381–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic). See [121].

Bencherif-Madani:2007:HSP

[159] Abdellatif Benchérif-Madani and Étienne Pardoux. Homogenization of a semilinear parabolic PDE with locally periodic coefficients: a probabilistic approach. ESAIM: Probability and Statistics, 11 (??):385–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pinelis:2007:TBC

[160] Iosif Pinelis. Toward the best constant factor for the Rademacher–Gaussian tail comparison. ESAIM: Probability and Statistics, 11(??):412–??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Douc:2007:MVI

[161] R. Douc, A. Guillin, J.-M. Marin, and C. P. Robert. Minimum variance importance sampling via Population Monte Carlo. ESAIM: Probability and Statistics, 11(??):427–??, February 2007. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

El-Nouty:2007:FMF

[162] Charles El-Nouty. The fractional mixed fractional Brownian motion and fractional Brownian sheet. ESAIM:

Probability and Statistics, 11(??):448– ??, February 2007. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Berkaoui:2008:ESS

[163] Abdel Berkaoui, Mireille Bossy, and Awa Diop. Euler scheme for SDEs with non-Lipschitz diffusion coefficient: strong convergence. ESAIM: Probability and Statistics, 12(??):1–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cattiaux:2008:DBA

[164] Patrick Cattiaux and Arnaud Guillin.

Deviation bounds for additive functionals of Markov processes. ESAIM: Probability and Statistics, 12(??):12–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bierme:2008:EAG

[165] Hermine Biermé and Frédéric Richard.

Estimation of anisotropic Gaussian fields through radon transform. ESAIM: Probability and Statistics, 12(??):30–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Miao:2008:CIS

[166] Yu Miao. Concentration inequalities for semi-bounded martingales. ESAIM: Probability and Statistics, 12(??):51–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Roynette:2008:ABH

[167] Bernard Roynette, Pierre Vallois, and Agnés Volpi. Asymptotic behavior of the hitting time, overshoot and undershoot for some Lévy processes. ESAIM: Probability and Statistics, 12(??):58–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kleptsyna:2008:SPF

[168] Marina L. Kleptsyna, Alain Le Breton, and Michel Viot. Separation principle in the fractional Gaussian linear-quadratic regulator problem with partial observation. ESAIM: Probability and Statistics, 12(??):94–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Graf:2008:DMQ

[169] Siegfried Graf, Harald Luschgy, and Gilles Pagés. Distortion mismatch in the quantization of probability measures. ESAIM: Probability and Statistics, 12(??):127–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bardet:2008:DLC

[170] Jean-Marc Bardet, Paul Doukhan, Gabriel Lang, and Nicolas Ragache. Dependent Lindeberg central limit theorem and some applications. ESAIM: Probability and Statistics, 12(??):154–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lavielle:2008:RTL

[171] Marc Lavielle and Carenne Ludeña. Random thresholds for linear model selection. ESAIM: Probability and Statistics, 12(??):173–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Donnet:2008:PIM

[172] Sophie Donnet and Adeline Samson. Parametric inference for mixed models defined by stochastic differential equations. ESAIM: Probability and Statistics, 12(??):196–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Galves:2008:EIV

[173] Antonio Galves, Véronique Maume-Deschamps, and Bernard Schmitt. Exponential inequalities for VLMC empirical trees. ESAIM: Probability and Statistics, 12(??):219–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tudor:2008:ARP

[174] Ciprian A. Tudor. Analysis of the Rosenblatt process. ESAIM: Probability and Statistics, 12(??):230–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Zitt:2008:FIU

[175] Pierre-André Zitt. Functional inequalities and uniqueness of the Gibbs measure — from log-Sobolev to Poincaré. ESAIM: Probability and Statistics, 12 (??):258–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Menneteau:2008:MLT

[176] Ludovic Menneteau. Multidimensional limit theorems for smoothed extreme value estimates of point processes boundaries. ESAIM: Probability and Statistics, 12(??):273–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chretien:2008:EAT

[177] Stéphane Chrétien and Alfred O. Hero.

On EM algorithms and their proximal generalizations. ESAIM: Probability and Statistics, 12(??):308–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ermakov:2008:MBE

[178] Mikhail Ermakov. Minimax and Bayes estimation in deconvolution problem. ESAIM: Probability and Statistics, 12 (??):327–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tran:2008:LPL

[179] Viet Chi Tran. Large population limit and time behaviour of a stochastic particle model describing an age-structured population. ESAIM: Probability and Statistics, 12(??):345–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lejay:2008:SDE

[180] Antoine Lejay. Stochastic differential equations driven by processes generated by divergence form operators II: convergence results. ESAIM: Probability and Statistics, 12(??):387–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Imkeller:2008:MBS

[181] Peter Imkeller and Ilya Pavlyukevich. Metastable behaviour of small noise Lévy-driven diffusions. ESAIM: Probability and Statistics, 12(??):412–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Alquier:2008:DEQ

[182] Pierre Alquier. Density estimation with quadratic loss: a confidence intervals method. ESAIM: Probability and Statistics, 12(??):438–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chaabane:2008:TLA

[183] Faouzi Chaabane and Ahmed Kebaier.

Théorémes limites avec poids pour les martingales vectorielles à temps continu. (French) [Weighted limit theorems for continuous-time vectorial martingales]. ESAIM: Probability and Statistics, 12 (??):464–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Collet:2008:LSI

[184] Jean-François Collet and Florent Malrieu. Logarithmic Sobolev inequalities for inhomogeneous Markov semigroups. ESAIM: Probability and Statistics, 12 (??):492–??, April 2008. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hengartner:2009:AUD

[185] Nicolas W. Hengartner and Éric Matzner-Løber. Asymptotic unbiased density estimators. ESAIM: Probability and Statistics, 13(??):1–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cenac:2009:DST

[186] Peggy Cénac, Brigitte Chauvin, Stéphane Ginouillac, and Nicolas Pouyanne. Digital search trees and chaos game representation. ESAIM: Probability and Statistics, 13(??):15–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chambaz:2009:NHS

[187] Antoine Chambaz and Catherine Matias. Number of hidden states and memory: a joint order estimation problem for Markov chains with Markov regime. ESAIM: Probability and Statistics, 13 (??):38–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boivin:2009:TEH

[188] Daniel Boivin. Tail estimates for homogenization theorems in random media. ESAIM: Probability and Statistics, 13(??):51–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sauve:2009:HSN

[189] Marie Sauvé. Histogram selection in non Gaussian regression. ESAIM: Probability and Statistics, 13(??):70–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Martin-Magniette:2009:EHF

[190] Marie-Laure Martin-Magniette and Marie-Luce Taupin. Estimation of the hazard function in a semiparametric model with covariate measurement error. ESAIM: Probability and Statistics, 13(??):87–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Godet:2009:LPL

[191] Fanny Godet. Linear prediction of longrange dependent time series. ESAIM: Probability and Statistics, 13(??):115–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Schick:2009:PEH

[192] Anton Schick and Wolfgang Wefelmeyer.

Plug-in estimators for higher-order transition densities in autoregression. ESAIM: Probability and Statistics, 13 (??):135–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Roynette:2009:PMB

[193] Bernard Roynette, Pierre Vallois, and Marc Yor. Penalisations of multidimensional Brownian motion, VI. ESAIM: Probability and Statistics, 13(??):152–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Davies:2009:CAH

[194] Laurie Davies, Ursula Gather, Dan Nordman, and Henrike Weinert. A comparison of automatic histogram constructions. ESAIM: Probability and Statistics, 13(??):181–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chaleyat-Maurel:2009:FWF

[195] Mireille Chaleyat-Maurel and Valentine Genon-Catalot. Filtering the Wright– Fisher diffusion. ESAIM: Probability and Statistics, 13(??):197–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sagna:2009:URO

[196] Abass Sagna. Universal L^{s}-rate-optimality of L^{r}-optimal quantizers by dilatation and contraction. ESAIM: Probability and Statistics, 13(??):218–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coutin:2009:EGP

[197] Laure Coutin and Nicolas Victoir. Enhanced Gaussian processes and applications. ESAIM: Probability and Statistics, 13(??):247–??, January 2009. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rackauskas:2009:HIP

[198] Alfredas Račkauskas and Charles Suquet. Hölderian invariance principle for Hilbertian linear processes. ESAIM: Probability and Statistics, 13(??):261–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pons:2009:ETF

[199] Odile Pons. Estimation and tests in finite mixture models of nonparametric densities. ESAIM: Probability and Statistics, 13(??):276–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Withers:2009:PCG

[200] Christopher S. Withers and Saralees Nadarajah. Power of a class of goodnessof-fit tests I. ESAIM: Probability and Statistics, 13(??):283–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Azais:2009:LRT

[201] Jean-Marc Azaïs, Élisabeth Gassiat, and Cécile Mercadier. The likelihood ratio test for general mixture models with or without structural parameter. ESAIM: Probability and Statistics, 13(??):301–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Karaliopoulou:2009:NWO

[202] Margarita Karaliopoulou. On the number of word occurrences in a semi-Markov sequence of letters. ESAIM: Probability and Statistics, 13(??):328–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lowe:2009:CBC

[203] Matthias Löwe and Franck Vermet. Capacity bounds for the CDMA system and a neural network: a moderate deviations approach. ESAIM: Probability and Statistics, 13(??):343–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kahn:2009:MSQ

[204] Jonas Kahn. Model selection for quantum homodyne tomography. ESAIM: Probability and Statistics, 13(??):363–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Barthe:2009:MVL

[205] Franck Barthe and Neil O’Connell.

Matchings and the variance of Lipschitz functions. ESAIM: Probability and Statistics, 13(??):400–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kabluchko:2009:STM

[206] Zakhar Kabluchko and Axel Munk.

Shao’s theorem on the maximum of standardized random walk increments for multidimensional arrays. ESAIM: Probability and Statistics, 13(??):409–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pene:2009:TRW

[207] Françoise Péne. Transient random walk in Z^{2} with stationary orientations. ESAIM: Probability and Statistics, 13 (??):417–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Akhavi:2009:RRB

[208] Ali Akhavi, Jean-François Marckert, and Alain Rouault. On the reduction of a random basis. ESAIM: Probability and Statistics, 13(??):437–??, January 2009. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Eichelsbacher:2010:PLM

[209] Peter Eichelsbacher and Tomasz Schreiber. Process level moderate deviations for stabilizing functionals. ESAIM: Probability and Statistics, 14(??):1–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bailleul:2010:WDR

[210] Ismael Bailleul and Albert Raugi. Where does randomness lead in spacetime? ESAIM: Probability and Statistics, 14 (??):16–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Barbour:2010:CBP

[211] Andrew D. Barbour. Coupling a branching process to an infinite dimensional epidemic process. ESAIM: Probability and Statistics, 14(??):53–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Roynette:2010:LLT

[212] Bernard Roynette and Marc Yor. Local limit theorems for Brownian additive functionals and penalisation of Brownian paths, IX. ESAIM: Probability and Statistics, 14(??):65–??, ???? 2010. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Luschgy:2010:AOQ

[213] Harald Luschgy, Gilles Pagés, and Benedikt Wilbertz. Asymptotically optimal quantization schemes for Gaussian processes on Hilbert spaces. ESAIM: Probability and Statistics, 14(??):93–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Miclo:2010:ATD

[214] Laurent Miclo. On absorption times and Dirichlet eigenvalues. ESAIM: Probability and Statistics, 14(??):117–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gannaz:2010:ADE

[215] Iréne Gannaz and Olivier Wintenberger.

Adaptive density estimation under weak dependence. ESAIM: Probability and Statistics, 14(??):151–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Loubes:2010:PEN

[216] Jean-Michel Loubes and Carenne Ludeña. Penalized estimators for non linear inverse problems. ESAIM: Probability and Statistics, 14(??):173–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bolley:2010:QCI

[217] François Bolley. Quantitative concentration inequalities on sample path space for mean field interaction. ESAIM: Probability and Statistics, 14(??):192–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Avram:2010:STL

[218] Florin Avram, Nikolai Leonenko, and Ludmila Sakhno. On a Szegő type limit theorem, the Hölder–Young–Brascamp– Lieb inequality, and the asymptotic theory of integrals and quadratic forms of stationary fields. ESAIM: Probability and Statistics, 14(??):210–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Downarowicz:2010:SCT

[219] T. Downarowicz, Y. Lacroix, and D. Léandri. Spontaneous clustering in theoretical and some empirical stationary processes. ESAIM: Probability and Statistics, 14(??):256–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Carmona:2010:DPR

[220] Philippe Carmona. Directed polymer in random environment and last passage percolation. ESAIM: Probability and Statistics, 14(??):263–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Grigorescu:2010:SSS

[221] Ilie Grigorescu and Min Kang. Steady state and scaling limit for a traffic congestion model. ESAIM: Probability and Statistics, 14(??):271–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fekete:2010:BRW

[222] Eric Fekete. Branching random walks on binary search trees: convergence of the occupation measure. ESAIM: Probability and Statistics, 14(??):286–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guillotin-Plantard:2010:CLT

[223] Nadine Guillotin-Plantard and Clémentine Prieur. Central limit theorem for sampled sums of dependent random variables. ESAIM: Probability and Statistics, 14(??):299–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bartlett:2010:OSB

[224] Peter L. Bartlett, Shahar Mendelson, and Petra Philips. On the optimality of sample-based estimates of the expectation of the empirical minimizer. ESAIM: Probability and Statistics, 14(??):315–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gonchigdanzan:2010:ASF

[225] Khurelbaatar Gonchigdanzan. Almost sure functional limit theorem for the product of partial sums. ESAIM: Probability and Statistics, 14(??):338–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Leonard:2010:EPD

[226] Christian Léonard. Entropic projections and dominating points. ESAIM: Probability and Statistics, 14(??):343–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Allassonniere:2010:SAB

[227] Stéphanie Allassonniére and Estelle Kuhn. Stochastic algorithm for Bayesian mixture effect template estimation. ESAIM: Probability and Statistics, 14 (??):382–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marteau:2010:RHM

[228] Clément Marteau. Risk hull method for spectral regularization in linear statistical inverse problems. ESAIM: Probability and Statistics, 14(??):409–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pudlo:2010:LDF

[229] Pierre Pudlo. Large deviations and full Edgeworth expansions for finite Markov chains with applications to the analysis of genomic sequences. ESAIM: Probability and Statistics, 14(??):435–??, ???? 2010. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hoffmann:2011:PSH

[230] Marc Hoffmann and Damien Lamberton. Preface — supplement: In honor of Marc Yor. ESAIM: Probability and Statistics, 15(??):S1, ???? 2011. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Biane:2011:ESG

[231] Philippe Biane. Entrelacements de semigroupes provenant de paires de Gelfand. (French) [semi-group interlacings from Gelfand pairs]. ESAIM: Probability and Statistics, 15(??):S2, ???? 2011. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cox:2011:THD

[232] Alexander M. G. Cox, David Hobson, and Jan Obłój. Time-homogeneous diffusions with a given marginal at a random time. ESAIM: Probability and Statistics, 15(??):S11, ???? 2011. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Follmer:2011:LMF

[233] Hans Föllmer and Philip Protter. Local martingales and filtration shrinkage. ESAIM: Probability and Statistics, 15 (??):S25, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Najnudel:2011:NKA

[234] Joseph Najnudel and Ashkan Nikeghbali. A new kind of augmentation of filtrations. ESAIM: Probability and Statistics, 15(??):S39, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Peccati:2011:HSS

[235] Giovanni Peccati and Jean-Renaud Pycke. Hoeffding spaces and Specht modules. ESAIM: Probability and Statistics, 15(??):S58, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yano:2011:WIC

[236] Kouji Yano. Wiener integral for the coordinate process under the σ-finite measure unifying Brownian penalisations. ESAIM: Probability and Statistics, 15 (??):S69, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Akakpo:2011:EDD

[237] Nathalie Akakpo. Estimating a discrete distribution via histogram selection. ESAIM: Probability and Statistics, 15(??):1–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Alheety:2011:NSR

[238] Mustafa Ismaeel Alheety. A new stochastic restricted biased estimator under heteroscedastic or correlated error. ESAIM: Probability and Statistics, 15(??):30–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Maugis:2011:NAP

[239] Cathy Maugis and Bertrand Michel. A non asymptotic penalized criterion for Gaussian mixture model selection. ESAIM: Probability and Statistics, 15 (??):41–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Loubes:2011:LAN

[240] Jean-Michel Loubes and Davy Paindaveine. Local asymptotic normality property for lacunar wavelet series multifractal model. ESAIM: Probability and Statistics, 15(??):69–??, ???? 2011. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Duffy:2011:LDC

[241] Ken R. Duffy, Claudio Macci, and Giovanni Luca Torrisi. On the large deviations of a class of modulated additive processes. ESAIM: Probability and Statistics, 15(??):83–??, ???? 2011. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Balan:2011:TSH

[242] Raluca M. Balan. L_{p}-theory for the stochastic heat equation with infinitedimensional fractional noise. ESAIM: Probability and Statistics, 15(??):110–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Withers:2011:EDE

[243] Christopher S. Withers and Saralees Nadarajah. Expansions for the distribution of M-estimates with applications to the multi-tone problem. ESAIM: Probability and Statistics, 15(??):139–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lai:2011:CTM

[244] Tze Leng Lai, Qi-Man Shao, and Qiying Wang. Cramér type moderate deviations for Studentized U-statistics. ESAIM: Probability and Statistics, 15(??):168–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Privault:2011:SSG

[245] Nicolas Privault and Anthony Réveillac.

SURE shrinkage of Gaussian paths and signal identification. ESAIM: Probability and Statistics, 15(??):180–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Locherbach:2011:PND

[246] Eva Löcherbach, Dasha Loukianova, and Oleg Loukianov. Penalized nonparametric drift estimation for a continuously observed one-dimensional diffusion process. ESAIM: Probability and Statistics, 15(??):197–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ibrahim:2011:LDD

[247] Jean-Paul Ibrahim. Large deviations for directed percolation on a thin rectangle. ESAIM: Probability and Statistics, 15(??):217–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fournier:2011:SAL

[248] Nicolas Fournier. Simulation and approximation of Lévy-driven stochastic differential equations. ESAIM: Probability and Statistics, 15(??):233–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coiffard:2011:RFG

[249] Claire Coiffard. Random fractals generated by a local Gaussian process indexed by a class of functions. ESAIM: Probability and Statistics, 15(??):249–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Uchida:2011:EME

[250] Masayuki Uchida and Nakahiro Yoshida. Estimation for misspecified ergodic diffusion processes from discrete observations. ESAIM: Probability and Statistics, 15(??):270–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yurachkivsky:2011:LTM

[251] Andriy Yurachkivsky. Limit theorems for measure-valued processes of the level-exceedance type. ESAIM: Probability and Statistics, 15(??):291–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Maugis:2011:DDP

[252] Cathy Maugis and Bertrand Michel.

Data-driven penalty calibration: A case study for Gaussian mixture model selection. ESAIM: Probability and Statistics, 15(??):320–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Withers:2011:ERI

[253] Christopher S. Withers and Saralees Nadarajah. Expansions for repeated integrals of products with applications to the multivariate normal. ESAIM: Probability and Statistics, 15(??):340–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rhodes:2011:KFL

[254] Rémi Rhodes and Vincent Vargas. KPZ formula for log-infinitely divisible multifractal random measures. ESAIM: Probability and Statistics, 15(??):358–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Voisin:2011:DMF

[255] Guillaume Voisin. Dislocation measure of the fragmentation of a general Lévy tree. ESAIM: Probability and Statistics, 15(??):372–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gallesco:2011:NSW

[256] Christophe Gallesco, Sebastian Müller, and Serguei Popov. A note on spider walks. ESAIM: Probability and Statistics, 15(??):390–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Aletti:2011:IDS

[257] Giacomo Aletti, Enea G. Bongiorno, and Vincenzo Capasso. Integration in a dynamical stochastic geometric framework. ESAIM: Probability and Statistics, 15 (??):402–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Penisson:2011:CTM

[258] Sophie Pénisson. Continuous-time multitype branching processes conditioned on very late extinction. ESAIM: Probability and Statistics, 15(??):417–??, ???? 2011. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Istas:2012:ESS

[259] Jacques Istas and Serge Cohen. Editorial — Spring School Mons: Random differential equations and Gaussian fields. ESAIM: Probability and Statistics, 16 (??):Sp–1, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cohen:2012:SGR

[260] S. Cohen and M. A. Lifshits. Stationary Gaussian random fields on hyperbolic spaces and on Euclidean spheres. ESAIM: Probability and Statistics, 16 (??):165–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Istas:2012:MIF

[261] Jacques Istas. Manifold indexed fractional fields. ESAIM: Probability and Statistics, 16(??):222–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baudoin:2012:STE

[262] Fabrice Baudoin. Stochastic Taylor expansions and heat kernel asymptotics. ESAIM: Probability and Statistics, 16 (??):453–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coutin:2012:RPS

[263] Laure Coutin. Rough paths via sewing lemma. ESAIM: Probability and Statistics, 16(??):479–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nane:2012:TFB

[264] Erkan Nane, Dongsheng Wu, and Yimin Xiao. α-time fractional Brownian motion: PDE connections and local times. ESAIM: Probability and Statistics, 16 (??):1–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Olteanu:2012:APA

[265] Madalina Olteanu and Joseph Rynkiewicz.

Asymptotic properties of autoregressive regime-switching models. ESAIM: Probability and Statistics, 16(??):25–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hirsch:2012:NPK

[266] Francis Hirsch and Bernard Roynette.

A new proof of Kellerer’s theorem. ESAIM: Probability and Statistics, 16 (??):48–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lerasle:2012:ANA

[267] Matthieu Lerasle. Adaptive nonasymptotic confidence balls in density estimation. ESAIM: Probability and Statistics, 16(??):61–??, ???? 2012. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pinelis:2012:EDC

[268] Iosif Pinelis. Exponential deficiency of convolutions of densities. ESAIM: Probability and Statistics, 16(??):86–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Worms:2012:ESO

[269] Julien Worms and Rym Worms. Estimation of second order parameters using probability weighted moments. ESAIM: Probability and Statistics, 16(??):97–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Doku-Amponsah:2012:AEP

[270] Kwabena Doku-Amponsah. Asymptotic equipartition properties for simple hierarchical and networked structures. ESAIM: Probability and Statistics, 16 (??):114–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nkurunziza:2012:SSS

[271] Sévérien Nkurunziza. Shrinkage strategies in some multiple multi-factor dynamical systems. ESAIM: Probability and Statistics, 16(??):139–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Favetto:2012:AVC

[272] Benjamin Favetto. On the asymptotic variance in the central limit theorem for particle filters. ESAIM: Probability and Statistics, 16(??):151–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Herrmann:2012:SSP

[273] Samuel Herrmann and Julian Tugaut.

Self-stabilizing processes: uniqueness problem for stationary measures and convergence rate in the small-noise limit. ESAIM: Probability and Statistics, 16 (??):277–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Beiglbock:2012:GDM

[274] Mathias Beiglböck, Christian Léonard, and Walter Schachermayer. A generalized dual maximizer for the Monge–Kantorovich transport problem. ESAIM: Probability and Statistics, 16 (??):306–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lachal:2012:STB

[275] Aimé Lachal. Sojourn time in Z^{+} for the Bernoulli random walk on Z. ESAIM: Probability and Statistics, 16(??):324–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clausel:2012:LFB

[276] Marianne Clausel. Lacunary fractional Brownian motion. ESAIM: Probability and Statistics, 16(??):352–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Laurent:2012:BFR

[277] Stéphane Laurent and Catherine Legrand. A Bayesian framework for the ratio of two Poisson rates in the context of vaccine efficacy trials. ESAIM: Probability and Statistics, 16(??):375–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Malric:2012:DPI

[278] Marc Malric. Density of paths of iterated Lévy transforms of Brownian motion. ESAIM: Probability and Statistics, 16(??):399–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hu:2012:TAS

[279] Zhishui Hu, John Robinson, and Qiying Wang. Tail approximations for samples from a finite population with applications to permutation tests. ESAIM: Probability and Statistics, 16(??):425–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bubeck:2012:HIA

[280] Sébastien Bubeck, Marina Meilă, and Ulrike von Luxburg. How the initialization affects the stability of the κmeans algorithm. ESAIM: Probability and Statistics, 16(??):436–??, ???? 2012. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Aubry:2013:LDQ

[281] Jean-Marie Aubry and Marguerite Zani. Large deviations for quasiarithmetically self-normalized random variables. ESAIM: Probability and Statistics, 17(??):1–??, ???? 2013. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kawai:2013:LAN

[282] Reiichiro Kawai and Hiroki Masuda.

Local asymptotic normality for normal inverse Gaussian Lévy processes with high-frequency sampling. ESAIM: Probability and Statistics, 17(??):13–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Schmisser:2013:NED

[283] Emeline Schmisser. Nonparametric estimation of the derivatives of the stationary density for stationary processes. ESAIM: Probability and Statistics, 17 (??):33–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rossignol:2013:LLD

[284] Raphaël Rossignol and Marie Théret.

Lower large deviations for the maximal flow through tilted cylinders in twodimensional first passage percolation. ESAIM: Probability and Statistics, 17 (??):70–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lelong:2013:ANR

[285] Jérôme Lelong. Asymptotic normality of randomly truncated stochastic algorithms. ESAIM: Probability and Statistics, 17(??):105–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Neumann:2013:CLT

[286] Michael H. Neumann. A central limit theorem for triangular arrays of weakly dependent random variables, with applications in statistics. ESAIM: Probability and Statistics, 17(??):120–??, ???? 2013.

CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

LeGuevel:2013:IMH

[287] Ronan Le Guével and Jacques Lévy Véhel. Incremental moments and hölder exponents of multifractional multistable processes. ESAIM: Probability and Statistics, 17(??):135–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cenac:2013:CMA

[288] Peggy Cénac. On the convergence of moments in the almost sure central limit theorem for stochastic approximation algorithms. ESAIM: Probability and Statistics, 17(??):179–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Locherbach:2013:PDB

[289] Eva Löcherbach and Dasha Loukianova. Polynomial deviation bounds for recurrent Harris processes having general state space. ESAIM: Probability and Statistics, 17(??):195–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Istas:2013:MBF

[290] Jacques Istas. Multifractional Brownian fields indexed by metric spaces with distances of negative type. ESAIM: Probability and Statistics, 17(??):219–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Withers:2013:FFE

[291] Christopher S. Withers and Saralees Nadarajah. Fixed-α and fixed-β efficiencies. ESAIM: Probability and Statistics, 17(??):224–??, ???? 2013. CODEN ????

ISSN 1292-8100 (print), 1262-3318 (electronic).

DiBernardino:2013:PEL

[292] Elena Di Bernardino, Thomas Laloë, Véronique Maume-Deschamps, and Clémentine Prieur. Plug-in estimation of level sets in a non-compact setting with applications in multivariate risk theory. ESAIM: Probability and Statistics, 17 (??):236–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gallesco:2013:MTI

[293] Christophe Gallesco. Meeting time of independent random walks in random environment. ESAIM: Probability and Statistics, 17(??):257–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Grinberg:2013:SDC

[294] Nastasiya F. Grinberg. Semimartingale decomposition of convex functions of continuous semimartingales by Brownian perturbation. ESAIM: Probability and Statistics, 17(??):293–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bertrand:2013:LEH

[295] Pierre Raphaël Bertrand, Mehdi Fhima, and Arnaud Guillin. Local estimation of the Hurst index of multifractional Brownian motion by increment ratio statistic method. ESAIM: Probability and Statistics, 17(??):307–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chagny:2013:PVG

[296] Gaëlle Chagny. Penalization versus Goldenshluger–Lepski strategies in warped bases regression. ESAIM: Probability and Statistics, 17(??):328–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cuevas:2013:TUC

[297] Antonio Cuevas, Ricardo Fraiman, and László Györfi. Towards a universally consistent estimator of the Minkowski content. ESAIM: Probability and Statistics, 17(??):359–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Maier:2013:HRG

[298] Markus Maier, Ulrike von Luxburg, and Matthias Hein. How the result of graph clustering methods depends on the construction of the graph. ESAIM: Probability and Statistics, 17(??):370–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chauveau:2013:SMH

[299] Didier Chauveau and Pierre Vandekerkhove. Smoothness of Metropolis– Hastings algorithm and application to entropy estimation. ESAIM: Probability and Statistics, 17(??):419–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dette:2013:GFT

[300] Holger Dette and Kemal Sen. Goodnessof-fit tests in long-range dependent processes under fixed alternatives. ESAIM: Probability and Statistics, 17(??):432–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hirsch:2013:VP

[301] Francis Hirsch and Bernard Roynette.

On R^{d}-valued peacocks. ESAIM: Probability and Statistics, 17(??):444–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fuchs:2013:MCM

[302] Florian Fuchs and Robert Stelzer. Mixing conditions for multivariate infinitely divisible processes with an application to mixed moving averages and the supOU stochastic volatility model. ESAIM: Probability and Statistics, 17(??):455–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fraysse:2013:WMP

[303] Aurelia Fraysse. Why minimax is not that pessimistic. ESAIM: Probability and Statistics, 17(??):472–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rochet:2013:AHT

[304] Paul Rochet. Adaptive hard-thresholding for linear inverse problems. ESAIM: Probability and Statistics, 17(??):485–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bercu:2013:SAA

[305] Bernard Bercu and Frédéric Proïa. A sharp analysis on the asymptotic behavior of the Durbin–Watson statistic for the first-order autoregressive process. ESAIM: Probability and Statistics, 17 (??):500–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Granovsky:2013:ACS

[306] Boris L. Granovsky. Asymptotics of counts of small components in random structures and models of coagulationfragmentation. ESAIM: Probability and Statistics, 17(??):531–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Callegaro:2013:CEF

[307] Giorgia Callegaro, Monique Jeanblanc, and Behnaz Zargari. Carthaginian enlargement of filtrations. ESAIM: Probability and Statistics, 17(??):550–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dombry:2013:MMH

[308] Clément Dombry and Ingemar Kaj. Moment measures of heavy-tailed renewal point processes: asymptotics and applications. ESAIM: Probability and Statistics, 17(??):567–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coeurjolly:2013:WAM

[309] Jean-François Coeurjolly, Pierre-Olivier Amblard, and Sophie Achard. Wavelet analysis of the multivariate fractional Brownian motion. ESAIM: Probability and Statistics, 17(??):592–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lefevre:2013:CPE

[310] Claude Lefévre and Sergey Utev. Convolution property and exponential bounds for symmetric monotone densities. ESAIM: Probability and Statistics, 17 (??):605–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Budhiraja:2013:DTM

[311] Amarjit Budhiraja, Pierre Del Moral, and Sylvain Rubenthaler. Discrete time Markovian agents interacting through a potential. ESAIM: Probability and Statistics, 17(??):614–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Charlier:2013:NSC

[312] Benjamin Charlier. Necessary and sufficient condition for the existence of a Fréchet mean on the circle. ESAIM: Probability and Statistics, 17(??):635–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Meynet:2013:OIL

[313] Caroline Meynet. An ℓ_{1}-oracle inequality for the lasso in finite mixture Gaussian regression models. ESAIM: Probability and Statistics, 17(??):650–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cohen:2013:PBC

[314] S. X. Cohen and E. Le Pennec.

Partition-based conditional density estimation. ESAIM: Probability and Statistics, 17(??):672–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Maugis-Rabusseau:2013:ADE

[315] C. Maugis-Rabusseau and B. Michel.

Adaptive density estimation for clustering with Gaussian mixtures. ESAIM: Probability and Statistics, 17(??):698–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Reichenbachs:2013:MDC

[316] Anselm Reichenbachs. Moderate deviations for a Curie–Weiss model with dynamical external field. ESAIM: Probability and Statistics, 17(??):725–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rachdi:2013:RBN

[317] Nabil Rachdi, Jean-Claude Fort, and Thierry Klein. Risk bounds for new Mestimation problems. ESAIM: Probability and Statistics, 17(??):740–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lang:2013:TRS

[318] Gabriel Lang and Eric Marcon. Testing randomness of spatial point patterns with the Ripley statistic. ESAIM: Probability and Statistics, 17(??):767–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yor:2013:OPL

[319] Marc Yor. Sur l’œuvre de Paul Lévy. (French) [On the work of Paul Lévy]. ESAIM: Probability and Statistics, 17 (??):789–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bertoin:2013:PLA

[320] Jean Bertoin. Paul Lévy et l’arithmétique des lois de probabilités. (French) [Paul Lévy and the arithmetic of the laws of probability]. ESAIM: Probability and Statistics, 17(??):790–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

LeGall:2013:PLM

[321] Jean-François Le Gall. Paul Lévy et le mouvement brownien. (French) [Paul Lévy and Brownian motion]. ESAIM: Probability and Statistics, 17(??):795–??, ???? 2013. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Antoniadis:2014:NRE

[322] Anestis Antoniadis, Marianna Pensky, and Theofanis Sapatinas. Nonparametric regression estimation based on spatially inhomogeneous data: minimax global convergence rates and adaptivity. ESAIM: Probability and Statistics, 18(??):1–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clausel:2014:WEL

[323] M. Clausel, F. Roueff, M. S. Taqqu, and C. Tudor. Wavelet estimation of the long memory parameter for Hermite polynomial of Gaussian processes. ESAIM: Probability and Statistics, 18(??):42–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gendre:2014:MSE

[324] Xavier Gendre. Model selection and estimation of a component in additive regression. ESAIM: Probability and Statistics, 18(??):77–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Theret:2014:ULD

[325] Marie Theret. Upper large deviations for maximal flows through a tilted cylinder. ESAIM: Probability and Statistics, 18(??):117–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dziedziul:2014:DSE

[326] Karol Dziedziul and Bogdan Ćmiel. Density smoothness estimation problem using a wavelet approach. ESAIM: Probability and Statistics, 18(??):130–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baumgarten:2014:SPA

[327] Christoph Baumgarten. Survival probabilities of autoregressive processes. ESAIM: Probability and Statistics, 18 (??):145–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Scholler:2014:TCD

[328] Julie Scholler. On the time constant in a dependent first passage percolation model. ESAIM: Probability and Statistics, 18(??):171–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Arnaudon:2014:MCM

[329] Marc Arnaudon and Laurent Miclo. Means in complete manifolds: uniqueness and approximation. ESAIM: Probability and Statistics, 18(??):185–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kovtun:2014:IMI

[330] Mikhail Kovtun, Igor Akushevich, and Anatoliy Yashin. On identifiability of mixtures of independent distribution laws. ESAIM: Probability and Statistics, 18(??):207–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Elie:2014:ACB

[331] Romuald Elie and Idris Kharroubi.

Adding constraints to BSDEs with jumps: an alternative to multidimensional reflections. ESAIM: Probability and Statistics, 18(??):233–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lescornel:2014:URE

[332] Héléne Lescornel, Jean-Michel Loubes, and Claudie Chabriac. Unbiased risk estimation method for covariance estimation. ESAIM: Probability and Statistics, 18(??):251–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sabbah:2014:UCB

[333] Camille Sabbah. Uniform confidence bands for local polynomial quantile estimators. ESAIM: Probability and Statistics, 18(??):265–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dedecker:2014:EAM

[334] Jérôme Dedecker, Adeline Samson, and Marie-Luce Taupin. Estimation in autoregressive model with measurement error. ESAIM: Probability and Statistics, 18(??):277–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Penda:2014:MDD

[335] S. Valére Bitseki Penda, Hacéne Djellout, and Frédéric Proïa. Moderate deviations for the Durbin–Watson statistic related to the first-order autoregressive process. ESAIM: Probability and Statistics, 18(??):308–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gugushvili:2014:CNP

[336] Shota Gugushvili and Peter Spreij. Consistent non-parametric Bayesian estimation for a time-inhomogeneous Brownian motion. ESAIM: Probability and Statistics, 18(??):332–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Janon:2014:ANE

[337] Alexandre Janon, Thierry Klein, Agnés Lagnoux, Maëlle Nodet, and Clémentine Prieur. Asymptotic normality and efficiency of two Sobol index estimators. ESAIM: Probability and Statistics, 18 (??):342–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

deSaporta:2014:RCB

[338] Benoîte de Saporta, Anne Gégout-Petit, and Laurence Marsalle. Random coefficients bifurcating autoregressive processes. ESAIM: Probability and Statistics, 18(??):365–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bertin:2014:AED

[339] Karine Bertin and Nicolas Klutchnikoff. Adaptive estimation of a density function using beta kernels. ESAIM: Probability and Statistics, 18(??):400–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Herbin:2014:ASL

[340] Erick Herbin, Benjamin Arras, and Geoffroy Barruel. From almost sure local regularity to almost sure Hausdorff dimension for Gaussian fields. ESAIM: Probability and Statistics, 18(??):418–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Villemonais:2014:GAM

[341] Denis Villemonais. General approximation method for the distribution of Markov processes conditioned not to be killed. ESAIM: Probability and Statistics, 18(??):441–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cattiaux:2014:LTS

[342] Patrick Cattiaux and Mawaki Manou-Abi. Limit theorems for some functionals with heavy tails of a discrete time Markov chain. ESAIM: Probability and Statistics, 18(??):468–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cornillon:2014:RBE

[343] Pierre-André Cornillon, N. W. Hengartner, and E. Matzner-Løber. Recursive bias estimation for multivariate regression smoothers. ESAIM: Probability and Statistics, 18(??):483–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Wang:2014:SAF

[344] Jian Wang. A simple approach to functional inequalities for non-local Dirichlet forms. ESAIM: Probability and Statistics, 18(??):503–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bogso:2014:AMT

[345] Antoine Marie Bogso. An application of multivariate total positivity to peacocks. ESAIM: Probability and Statistics, 18(??):514–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Genadot:2014:MPD

[346] A. Genadot and M. Thieullen. Multiscale Piecewise Deterministic Markov Process in infinite dimension: central limit theorem and Langevin approximation. ESAIM: Probability and Statistics, 18(??):541–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Longla:2014:DSC

[347] Martial Longla. On dependence structure of copula-based Markov chains. ESAIM: Probability and Statistics, 18 (??):570–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nguyen:2014:NED

[348] Van Hanh Nguyen and Catherine Matias. Nonparametric estimation of the density of the alternative hypothesis in a multiple testing setup. Application to local false discovery rate estimation. ESAIM: Probability and Statistics, 18 (??):584–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chassagneux:2014:DRB

[349] Jean-François Chassagneux and Stéphane Crépey. Doubly reflected BSDEs with call protection and their approximation. ESAIM: Probability and Statistics, 18 (??):613–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Girard:2014:USC

[350] Stéphane Girard, Armelle Guillou, and Gilles Stupfler. Uniform strong consistency of a frontier estimator using kernel regression on high order moments. ESAIM: Probability and Statistics, 18 (??):642–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mitov:2014:EAP

[351] Kosto V. Mitov and Saralees Nadarajah.

Extremal and additive processes generated by Pareto distributed random vectors. ESAIM: Probability and Statistics, 18(??):667–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Etore:2014:ESS

[352] Pierre Étoré and Miguel Martinez.

Exact simulation for solutions of one-dimensional Stochastic Differential Equations with discontinuous drift. ESAIM: Probability and Statistics, 18 (??):686–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hillion:2014:NDB

[353] Erwan Hillion, Oliver Johnson, and Yaming Yu. A natural derivative on [0, n] and a binomial Poincaré inequality. ESAIM: Probability and Statistics, 18(??):703–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kamatani:2014:LDM

[354] Kengo Kamatani. Local degeneracy of Markov chain Monte Carlo methods. ESAIM: Probability and Statistics, 18 (??):713–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Azais:2014:RNE

[355] Romain Azaïs. A recursive nonparametric estimator for the transition kernel of a piecewise-deterministic Markov process. ESAIM: Probability and Statistics, 18(??):726–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cleynen:2014:SPN

[356] Alice Cleynen and Emilie Lebarbier. Segmentation of the Poisson and negative binomial rate models: a penalized estimator. ESAIM: Probability and Statistics, 18(??):750–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sauve:2014:VST

[357] Marie Sauve and Christine Tuleau-Malot. Variable selection through CART. ESAIM: Probability and Statistics, 18(??):770–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marie:2014:GMR

[358] Nicolas Marie. A generalized meanreverting equation and applications. ESAIM: Probability and Statistics, 18 (??):799–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Song:2014:OSF

[359] Shiqi Song. Optional splitting formula in a progressively enlarged filtration. ESAIM: Probability and Statistics, 18 (??):829–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marckert:2014:CCS

[360] Jean-François Marckert and David Renault. Compact convex sets of the plane and probability theory. ESAIM: Probability and Statistics, 18(??):854–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Benhenni:2014:LPE

[361] Karim Benhenni and David Degras. Local polynomial estimation of the mean function and its derivatives based on functional data and regular designs. ESAIM: Probability and Statistics, 18 (??):881–??, ???? 2014. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guyader:2015:IIR

[362] Arnaud Guyader, Nick Hengartner, Nicolas Jégou, and Eric Matzner-Løber. Iterative isotonic regression. ESAIM: Probability and Statistics, 19(??):1–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pinelis:2015:EBC

[363] Iosif Pinelis. Exact bounds on the closeness between the Student and standard normal distributions. ESAIM: Probability and Statistics, 19(??):24–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bertail:2015:TIE

[364] Patrice Bertail, Emilie Chautru, and Stéphan Clémençon. Tail index estimation based on survey data. ESAIM: Probability and Statistics, 19(??):28–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fort:2015:CLT

[365] Gersende Fort. Central limit theorems for stochastic approximation with controlled Markov chain dynamics. ESAIM: Probability and Statistics, 19(??):60–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marquez-Carreras:2015:SSP

[366] David Márquez-Carreras. Small stochastic perturbations in a general fractional kinetic equation. ESAIM: Probability and Statistics, 19(??):81–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Barczy:2015:OSW

[367] Mátyás Barczy, Peter Kern, and Vincent Krause. Operator scaled Wiener bridges. ESAIM: Probability and Statistics, 19(??):100–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Butucea:2015:SVS

[368] Cristina Butucea, Yuri I. Ingster, and Irina A. Suslina. Sharp variable selection of a sparse submatrix in a highdimensional noisy matrix. ESAIM: Probability and Statistics, 19(??):115–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ourbih-Tari:2015:CMS

[369] Megdouda Ourbih-Tari and Sofia Guebli. A comparison of methods for selecting values of simulation input variables. ESAIM: Probability and Statistics, 19(??):135–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Profeta:2015:SLL

[370] Christophe Profeta. Some limiting laws associated with the integrated Brownian motion. ESAIM: Probability and Statistics, 19(??):148–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Vi:2015:HTM

[371] Le Vi and Etienne Pardoux. Height and the total mass of the forest of genealogical trees of a large population with general competition. ESAIM: Probability and Statistics, 19(??):172–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Karouf:2015:RBD

[372] Monia Karouf. Reflected BSDE’s with discontinuous barrier and time delayed generators. ESAIM: Probability and Statistics, 19(??):194–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sart:2015:MSP

[373] Mathieu Sart. Model selection for Poisson processes with covariates. ESAIM: Probability and Statistics, 19(??):204–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hirsch:2015:APC

[374] Christian Hirsch, Gerd Gaiselmann, and Volker Schmidt. Asymptotic properties of collective-rearrangement algorithms. ESAIM: Probability and Statistics, 19 (??):236–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cadre:2015:NRE

[375] Benoît Cadre and Lionel Truquet. Nonparametric regression estimation onto a Poisson point process covariate. ESAIM: Probability and Statistics, 19(??):251–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ailliot:2015:CML

[376] Pierre Ailliot and Françoise Pène. Consistency of the maximum likelihood estimate for non-homogeneous Markov switching models. ESAIM: Probability and Statistics, 19(??):268–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kovchegov:2015:OPS

[377] Yevgeniy Kovchegov and Neşe Yíldíz. Orthogonal polynomials for seminonparametric instrumental variables model. ESAIM: Probability and Statistics, 19(??):293–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bansaye:2015:WLL

[378] Vincent Bansaye and Chunmao Huang. Weak law of large numbers for some Markov chains along non homogeneous genealogies. ESAIM: Probability and Statistics, 19(??):307–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Allez:2015:CSE

[379] Romain Allez, Rémi Rhodes, and Vincent Vargas. Convergence of the spectrum of empirical covariance matrices for independent MRW processes. ESAIM: Probability and Statistics, 19 (??):327–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Brehier:2015:AAM

[380] Charles-Edouard Bréhier, Tony Lelièvre, and Mathias Rousset. Analysis of adaptive multilevel splitting algorithms in an idealized case. ESAIM: Probability and Statistics, 19(??):361–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Giuliano:2015:ARW

[381] Rita Giuliano and Claudio Macci.

Asymptotic results for weighted means of random variables which converge to a Dickman distribution, and some number theoretical applications. ESAIM: Probability and Statistics, 19(??):395–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Russo:2015:GNG

[382] Francesco Russo and Frederi Viens.

Gaussian and non-Gaussian processes of zero power variation. ESAIM: Probability and Statistics, 19(??):414–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Adamczak:2015:ECI

[383] Radosław Adamczak and Witold Bednorz. Exponential concentration inequalities for additive functionals of Markov chains. ESAIM: Probability and Statistics, 19(??):440–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bouguet:2015:QSC

[384] Florian Bouguet. Quantitative speeds of convergence for exposure to food contaminants. ESAIM: Probability and Statistics, 19(??):482–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Angst:2015:PBR

[385] Jürgen Angst. Poisson boundary of a relativistic diffusion in curved space-times: an example. ESAIM: Probability and Statistics, 19(??):502–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marie:2015:SRP

[386] Nicolas Marie. Sensitivities via rough paths. ESAIM: Probability and Statistics, 19(??):515–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fathi:2015:MLS

[387] Max Fathi. Modified logarithmic Sobolev inequalities for canonical ensembles. ESAIM: Probability and Statistics, 19(??):544–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mariucci:2015:AEI

[388] Ester Mariucci. Asymptotic equivalence for inhomogeneous jump diffusion processes and white noise. ESAIM: Probability and Statistics, 19(??):560–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rosenbaum:2015:SEF

[389] Mathieu Rosenbaum and Marc Yor.

Some explicit formulas for the Brownian bridge, Brownian meander and Bessel process under uniform sampling. ESAIM: Probability and Statistics, 19 (??):578–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boutsikas:2015:PGA

[390] Michael V. Boutsikas. Penultimate gamma approximation in the CLT for skewed distributions. ESAIM: Probability and Statistics, 19(??):590–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Andreoletti:2015:HMM

[391] Pierre Andreoletti, Dasha Loukianova, and Catherine Matias. Hidden Markov model for parameter estimation of a random walk in a Markov environment. ESAIM: Probability and Statistics, 19 (??):605–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ekstrom:2015:BST

[392] Erik Ekström and Juozas Vaicenavicius.

Bayesian sequential testing of the drift of a Brownian motion. ESAIM: Probability and Statistics, 19(??):626–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Devijver:2015:OIL

[393] Emilie Devijver. An ℓ_{1}-oracle inequality for the Lasso in multivariate finite mixture of multivariate Gaussian regression models. ESAIM: Probability and Statistics, 19(??):649–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Delattre:2015:EPP

[394] Maud Delattre, Valentine Genon-Catalot, and Adeline Samson. Estimation of population parameters in stochastic differential equations with random effects in the diffusion coefficient. ESAIM: Probability and Statistics, 19(??):671–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Penda:2015:DIB

[395] S. Valère Bitseki Penda. Deviation inequalities for bifurcating Markov chains on Galton–Watson tree. ESAIM: Probability and Statistics, 19(??):689–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

deCastro:2015:RPF

[396] Yohann de Castro and Alexandre Janon. Randomized pick-freeze for sparse Sobol indices estimation in high dimension. ESAIM: Probability and Statistics, 19 (??):725–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mazo:2015:WLS

[397] Gildas Mazo, Stéphane Girard, and Florence Forbes. Weighted least-squares inference for multivariate copulas based on dependence coefficients. ESAIM: Probability and Statistics, 19(??):746–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lifshits:2015:BBM

[398] Mikhail Lifshits and Ksenia Volkova. Bifractional Brownian motion: existence and border cases. ESAIM: Probability and Statistics, 19(??):766–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Breton:2015:IDF

[399] Jean-Christophe Breton and Renan Gobard. Infinite dimensional functional convergences in random balls model. ESAIM: Probability and Statistics, 19 (??):782–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Walter:2015:RES

[400] Clément Walter. Rare event simulation and splitting for discontinuous random variables. ESAIM: Probability and Statistics, 19(??):794–??, ???? 2015. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yoshida:2016:PLE

[401] Takuma Yoshida. Partially linear estimation using sufficient dimension reduction. ESAIM: Probability and Statistics, 20(??):1–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bonnefont:2016:NSG

[402] Michel Bonnefont, Aldéric Joulin, and Yutao Ma. A note on spectral gap and weighted Poincaré inequalities for some one-dimensional diffusions. ESAIM: Probability and Statistics, 20(??):18–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fischer:2016:CTM

[403] Markus Fischer and Giulia Livieri. Continuous time mean-variance portfolio optimization through the mean field approach. ESAIM: Probability and Statistics, 20(??):30–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Daly:2016:NDS

[404] Fraser Daly. Negative dependence and stochastic orderings. ESAIM: Probability and Statistics, 20(??):45–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dauxois:2016:ELC

[405] Jean-Yves Dauxois, Alexis Flesch, and Davit Varron. Empirical likelihood confidence bands for mean functions of recurrent events with competing risks and a terminal event. ESAIM: Probability and Statistics, 20(??):66–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nagy:2016:IDF

[406] Stanislav Nagy, Irène Gijbels, Marek Omelka, and Daniel Hlubinka. Integrated depth for functional data: statistical properties and consistency. ESAIM: Probability and Statistics, 20 (??):95–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coulson:2016:PAS

[407] Matthew Coulson, Robert E. Gaunt, and Gesine Reinert. Poisson approximation of subgraph counts in stochastic block models and a graphon model. ESAIM: Probability and Statistics, 20 (??):131–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gugushvili:2016:PCR

[408] Shota Gugushvili and Peter Spreij.

Posterior contraction rate for nonparametric Bayesian estimation of the dispersion coefficient of a stochastic differential equation. ESAIM: Probability and Statistics, 20(??):143–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sandric:2016:ELT

[409] Nikola Sandrić. Ergodicity of Lévytype processes. ESAIM: Probability and Statistics, 20(??):154–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Heyne:2016:MSC

[410] Gregor Heyne, Michael Kupper, Christoph Mainberger, and Ludovic Tangpi. Minimal supersolutions of convex BSDEs under constraints. ESAIM: Probability and Statistics, 20(??):178–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Krell:2016:SEJ

[411] Nathalie Krell. Statistical estimation of jump rates for a piecewise deterministic Markov processes with deterministic increasing motion and jump mechanism. ESAIM: Probability and Statistics, 20(??):196–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Benaych-Georges:2016:SAG

[412] Florent Benaych-Georges and Romain Couillet. Spectral analysis of the Gram matrix of mixture models. ESAIM: Probability and Statistics, 20(??):217–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Danka:2016:ABC

[413] Tivadar Danka and Gyula Pap. Asymptotic behavior of critical indecomposable multi-type branching processes with immigration. ESAIM: Probability and Statistics, 20(??):238–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hirsch:2016:MDS

[414] Christian Hirsch, David Neuhäuser, and Volker Schmidt. Moderate deviations for shortest-path lengths on random segment processes. ESAIM: Probability and Statistics, 20(??):261–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nourdin:2016:MLT

[415] Ivan Nourdin, Giovanni Peccati, Guillaume Poly, and Rosaria Simone. Multidimensional limit theorems for homogeneous sums: A survey and a general transfer principle. ESAIM: Probability and Statistics, 20(??):293–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kwemou:2016:NAO

[416] Marius Kwemou. Non-asymptotic oracle inequalities for the Lasso and Group Lasso in high dimensional logistic model. ESAIM: Probability and Statistics, 20 (??):309–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rodriguez-Casal:2016:FDD

[417] A. Rodríguez-Casal and P. Saavedra-Nieves. A fully data-driven method for estimating the shape of a point cloud. ESAIM: Probability and Statistics, 20 (??):332–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Albin:2016:ELT

[418] Patrik Albin, Enkelejd Hashorva, Lanpeng Ji, and Chengxiu Ling. Extremes and limit theorems for difference of chitype processes. ESAIM: Probability and Statistics, 20(??):349–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Doumas:2016:CCP

[419] Aristides V. Doumas and Vassilis G.

Papanicolaou. The coupon collector’s problem revisited: generalizing the double Dixie cup problem of Newman and Shepp. ESAIM: Probability and Statistics, 20(??):367–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cator:2016:CMT

[420] Eric Cator and Henk Don. Conditioned multi-type Galton–Watson trees. ESAIM: Probability and Statistics, 20 (??):400–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Privault:2016:PSC

[421] Nicolas Privault. Poisson sphere counting processes with random radii. ESAIM: Probability and Statistics, 20 (??):417–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sohl:2016:ACB

[422] Jakob Söhl and Mathias Trabs. Adaptive confidence bands for Markov chains and diffusions: Estimating the invariant measure and the drift. ESAIM: Probability and Statistics, 20(??):432–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fraiman:2016:SRI

[423] Ricardo Fraiman, Yanina Gimenez, and Marcela Svarc. Seeking relevant information from a statistical model. ESAIM: Probability and Statistics, 20 (??):463–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Popier:2016:LBB

[424] A. Popier. Limit behaviour of BSDE with jumps and with singular terminal condition. ESAIM: Probability and Statistics, 20(??):480–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boutahar:2016:TEM

[425] Mohamed Boutahar and Denys Pommeret. A test for the equality of monotone transformations of two random variables. ESAIM: Probability and Statistics, 20(??):510–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hopfner:2016:ELT

[426] Reinhard Höpfner, Eva Löcherbach, and Michèle Thieullen. Ergodicity and limit theorems for degenerate diffusions with time periodic drift. application to a stochastic Hodgkin–Huxley model. ESAIM: Probability and Statistics, 20 (??):527–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Barbu:2016:SSM

[427] Vlad Stefan Barbu, Guglielmo D’Amico, Raimondo Manca, and Filippo Petroni. Step semi-Markov models and application to manpower management. ESAIM: Probability and Statistics, 20(??):555–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Duboscq:2016:ASS

[428] Romain Duboscq and Renaud Marty.

Analysis of a splitting scheme for a class of random nonlinear partial differential equations. ESAIM: Probability and Statistics, 20(??):572–??, ???? 2016. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Drame:2017:NBB

[429] Ibrahima Drame, Etienne Pardoux, and A. B. Sow. Non-binary branching process and non-Markovian exploration process. ESAIM: Probability and Statistics, 21(??):1–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sart:2017:ECD

[430] Mathieu Sart. Estimating the conditional density by histogram type estimators and model selection. ESAIM: Probability and Statistics, 21(??):34–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Doumbia:2017:UMC

[431] Mahamadou Doumbia, Nadia Oudjane, and Xavier Warin. Unbiased Monte Carlo estimate of stochastic differential equations expectations. ESAIM: Probability and Statistics, 21(??):56–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Konakov:2017:SDP

[432] Valentin Konakov, Anna Kozhina, and Stéphane Menozzi. Stability of densities for perturbed diffusions and Markov chains. ESAIM: Probability and Statistics, 21(??):88–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Henry:2017:CLT

[433] Benoît Henry. Central limit theorem for supercritical binary homogeneous Crump–Mode–Jagers processes. ESAIM: Probability and Statistics, 21 (??):113–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cadre:2017:MRE

[434] Benoît Cadre, Nicolas Klutchnikoff, and Gaspar Massiot. Minimax regression estimation for Poisson coprocess. ESAIM: Probability and Statistics, 21(??):138–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chen:2017:QEA

[435] Jinwen Chen and Siqi Jian. Quasiergodicity for absorbing Markov processes via deviation inequality. ESAIM: Probability and Statistics, 21(??):159–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Owo:2017:PSB

[436] Jean-Marc Owo. L_{p}-solutions of backward doubly stochastic differential equations with stochastic Lipschitz condition and p ∈ (1, 2). ESAIM: Probability and Statistics, 21(??):168–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bernardino:2017:NUP

[437] Elena Di Bernardino and Didier Rullière. A note on upper-patched generators for Archimedean copulas. ESAIM: Probability and Statistics, 21 (??):183–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sang:2017:FRS

[438] Hailin Sang and Lin Ge. Further refinement of self-normalized Cramér-type moderate deviations. ESAIM: Probability and Statistics, 21(??):201–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Franceschi:2017:TIA

[439] S. Franceschi and Kilian Raschel.

Tutte’s invariant approach for Brownian motion reflected in the quadrant. ESAIM: Probability and Statistics, 21 (??):220–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Simoes:2017:PLD

[440] Jefferson Elbert Simões, Daniel R.

Figueiredo, and Valmir C. Barbosa. Power-law decay of the degree-sequence probabilities of multiple random graphs with application to graph isomorphism. ESAIM: Probability and Statistics, 21 (??):235–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Owhadi:2017:QRB

[441] Houman Owhadi and Clint Scovel.

Qualitative robustness in Bayesian inference. ESAIM: Probability and Statistics, 21(??):251–??, ???? 2017. CODEN ????

ISSN 1292-8100 (print), 1262-3318 (electronic).

Hoang:2017:EDK

[442] Van Ha Hoang. Estimating the division kernel of a size-structured population. ESAIM: Probability and Statistics, 21(??):275–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gaunt:2017:SCG

[443] Robert E. Gaunt. A Stein characterisation of the generalized hyperbolic distribution. ESAIM: Probability and Statistics, 21(??):303–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sarantsev:2017:TSI

[444] Andrey Sarantsev. Two-sided infinite systems of competing Brownian particles. ESAIM: Probability and Statistics, 21(??):317–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boyer:2017:RRW

[445] Jean-Baptiste Boyer. On the reflected random walk on R^{+}. ESAIM: Probability and Statistics, 21(??):350–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Antunes:2017:SLS

[446] Nelson Antunes, Vladas Pipiras, Patrice Abry, and Darryl Veitch. Small and large scale behavior of moments of Poisson cluster processes. ESAIM: Probability and Statistics, 21(??):369–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ciolek:2017:BPA

[447] Gabriela Ciołek and Paweł Potorski.

Bootstrapping periodically autoregressive models. ESAIM: Probability and Statistics, 21(??):394–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Navarro:2017:SHV

[448] Fabien Navarro and Adrien Saumard.

Slope heuristics and V-fold model selection in heteroscedastic regression using strongly localized bases. ESAIM: Probability and Statistics, 21(??):412–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Younso:2017:NCW

[449] Ahmad Younso. On nonparametric classification for weakly dependent functional processes. ESAIM: Probability and Statistics, 21(??):452–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Adamczak:2017:MEI

[450] Radosław Adamczak, Witold Bednorz, and Paweł Wolff. Moment estimates implied by modified log-Sobolev inequalities. ESAIM: Probability and Statistics, 21(??):467–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Debicki:2017:ETR

[451] Krzysztof Dȩbicki, Enkelejd Hashorva, and Peng Liu. Extremes of γ-reflected Gaussian processes with stationary increments. ESAIM: Probability and Statistics, 21(??):495–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cmiel:2017:VPE

[452] Bogdan Ćmiel and Teresa Ledwina. Validation of positive expectation dependence. ESAIM: Probability and Statistics, 21(??):536–??, ???? 2017. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Albin:2018:EVT

[453] Patrik Albin. On extreme value theory for group stationary Gaussian processes. ESAIM: Probability and Statistics, 22(??):1–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Newton:2018:MDD

[454] Nigel J. Newton. Manifolds of differentiable densities. ESAIM: Probability and Statistics, 22(??):19–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bigot:2018:CBW

[455] Jérémie Bigot and Thierry Klein. Characterization of barycenters in the Wasserstein space by averaging optimal transport maps. ESAIM: Probability and Statistics, 22(??):35–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clement:2018:AST

[456] Emmanuelle Clément, Arnaud Gloter, and Huong Nguyen. Asymptotics in small time for the density of a stochastic differential equation driven by a stable Lévy process. ESAIM: Probability and Statistics, 22(??):58–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Duroux:2018:IST

[457] Roxane Duroux and Erwan Scornet. Impact of subsampling and tree depth on random forests. ESAIM: Probability and Statistics, 22(??):96–??, ???? 2018. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Champagnat:2018:UCP

[458] Nicolas Champagnat and Denis Villemonais. Uniform convergence of penalized time-inhomogeneous Markov processes. ESAIM: Probability and Statistics, 22(??):129–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Beck:2018:CEO

[459] Nicholas Beck and Mélina Mailhot. A consistent estimator to the orthantbased tail value-at-risk. ESAIM: Probability and Statistics, 22(??):163–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kohatsu-Higa:2018:SFP

[460] Arturo Kohatsu-Higa and Gô Yûki.

Stochastic formulations of the parametrix method. ESAIM: Probability and Statistics, 22(??):178–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Brunel:2018:CPP

[461] Victor-Emmanuel Brunel. A changepoint problem and inference for segment signals. ESAIM: Probability and Statistics, 22(??):210–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Funke:2018:AND

[462] Benedikt Funke and Émeline Schmisser. Adaptive nonparametric drift estimation of an integrated jump diffusion process. ESAIM: Probability and Statistics, 22(??):236–??, ???? 2018. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Leon:2019:CLE

[463] José Rafael León, Luis-Ángel Rodríguez, and Roberto Ruggiero. Consistency of a likelihood estimator for stochastic damping Hamiltonian systems. Totally observed data. ESAIM: Probability and Statistics, 23(??):1–??, ???? 2019. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fan:2019:TFM

[464] Xiequan Fan and Jacques Lévy Véhel. Tempered fractional multistable motion and tempered multifractional stable motion. ESAIM: Probability and Statistics, 23(??):37–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bendikov:2019:RCC

[465] Alexander Bendikov and Wojciech Cygan. On the rate of convergence in the central limit theorem for hierarchical Laplacians. ESAIM: Probability and Statistics, 23(??):68–??, ???? 2019. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Letemplier:2019:LHS

[466] Julien Letemplier and Thomas Simon. On the law of homogeneous stable functionals. ESAIM: Probability and Statistics, 23(??):82–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Zhang:2019:MDC

[467] Xīlíng Zhāng. A multi-dimensional central limit bound and its application to the Euler approximation for Lévy– SDEs. ESAIM: Probability and Statistics, 23(??):112–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clement:2019:LPD

[468] Emmanuelle Clément, Arnaud Gloter, and Huong Nguyen. LAMN property for the drift and volatility parameters of a sde driven by a stable Lévy process. ESAIM: Probability and Statistics, 23(??):136–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Delattre:2019:MCT

[469] Sylvain Delattre and Nicolas Fournier.

On Monte-Carlo tree search for deterministic games with alternate moves and complete information. ESAIM: Probability and Statistics, 23(??):176–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chau:2019:FGR

[470] Huy N. Chau, Chaman Kumar, Miklós Rásonyi, and Sotirios Sabanis. On fixed gain recursive estimators with discontinuity in the parameters. ESAIM: Probability and Statistics, 23(??):217–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marque-Pucheu:2019:ESE

[471] Sophie Marque-Pucheu, Guillaume Perrin, and Josselin Garnier. Efficient sequential experimental design for surrogate modeling of nested codes. ESAIM: Probability and Statistics, 23(??):245–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mure:2019:OCB

[472] Joseph Muré. Optimal compromise between incompatible conditional probability distributions, with application to Objective Bayesian Kriging. ESAIM: Probability and Statistics, 23(??):271–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clemencon:2019:OSS

[473] Stephan Clémençon, Patrice Bertail, Emilie Chautru, and Guillaume Papa. Optimal survey schemes for stochastic gradient descent with applications to M-estimation. ESAIM: Probability and Statistics, 23(??):310–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gu:2019:SES

[474] Yu Gu. The 1D Schrödinger equation with a spacetime white noise: the average wave function. ESAIM: Probability and Statistics, 23(??):338–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bulinski:2019:SEC

[475] Alexander Bulinski and Alexey Kozhevin. Statistical estimation of conditional Shannon entropy. ESAIM: Probability and Statistics, 23(??):350–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cousin:2019:CSI

[476] A. Cousin, A. Janon, V. Maume-Deschamps, and I. Niang. On the consistency of Sobol indices with respect to stochastic ordering of model parameters. ESAIM: Probability and Statistics, 23(??):387–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Miclo:2019:CIQ

[477] Laurent Miclo. Complex intertwinings and quantification of discrete free motions. ESAIM: Probability and Statistics, 23(??):409–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dallaporta:2019:SRE

[478] Sandrine Dallaporta and Yohann De Castro. Sparse recovery from extreme eigenvalues deviation inequalities. ESAIM: Probability and Statistics, 23 (??):430–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lagnoux:2019:BRT

[479] Agnès Lagnoux, Thi Mong Ngoc Nguyen, and Frédéric Proïa. On the Bickel–Rosenblatt test of goodness-of-fit for the residuals of autoregressive processes. ESAIM: Probability and Statistics, 23(??):464–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Diehn:2019:MLE

[480] Manuel Diehn, Axel Munk, and Daniel Rudolf. Maximum likelihood estimation in hidden Markov models with inhomogeneous noise. ESAIM: Probability and Statistics, 23(??):492–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kroll:2019:ROE

[481] Martin Kroll. Rate optimal estimation of quadratic functionals in inverse problems with partially unknown operator and application to testing problems. ESAIM: Probability and Statistics, 23 (??):524–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yao:2019:BIM

[482] Nian Yao and Zhengliang Zhang. Beckner inequalities for Moebius measures on spheres. ESAIM: Probability and Statistics, 23(??):552–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lejay:2019:TCE

[483] Antoine Lejay, Ernesto Mordecki, and Soledad Torres. Two consistent estimators for the skew Brownian motion. ESAIM: Probability and Statistics, 23 (??):567–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Herve:2019:EGB

[484] Loïc Hervé, Sana Louhichi, and Françoise Pène. Exponential growth of branching processes in a general context of lifetimes and birthtimes dependence. ESAIM: Probability and Statistics, 23 (??):584–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Herve:2019:MEL

[485] Loïc Hervé, Sana Louhichi, and Françoise Pène. Multiplicative ergodicity of Laplace transforms for additive functional of Markov chains. ESAIM: Probability and Statistics, 23(??):607–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Marguet:2019:LLN

[486] Aline Marguet. A law of large numbers for branching Markov processes by the ergodicity of ancestral lineages. ESAIM: Probability and Statistics, 23(??):638–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Loffler:2019:WSH

[487] Matthias Löffler. Wald statistics in highdimensional PCA. ESAIM: Probability and Statistics, 23(??):662–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Xie:2019:TBC

[488] Junshan Xie and Gaoming Sun. A test for block circular symmetric covariance structure with divergent dimension. ESAIM: Probability and Statistics, 23 (??):672–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Inglot:2019:IEN

[489] Tadeusz Inglot, Teresa Ledwina, and Bogdan Ćmiel. Intermediate efficiency in nonparametric testing problems with an application to some weighted statistics. ESAIM: Probability and Statistics, 23(??):697–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lorek:2019:AMM

[490] Paweł Lorek. Antiduality and Möbius monotonicity: generalized coupon collector problem. ESAIM: Probability and Statistics, 23(??):739–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Duarte:2019:SCE

[491] Aline Duarte, Eva Löcherbach, and Guilherme Ost. Stability, convergence to equilibrium and simulation of nonlinear Hawkes processes with memory kernels given by the sum of Erlang kernels. ESAIM: Probability and Statistics, 23(??):770–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cerf:2019:GWB

[492] Raphaël Cerf and Joseba Dalmau. Galton–Watson and branching process representations of the normalized Perron–Frobenius eigenvector. ESAIM: Probability and Statistics, 23(??):797–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Nielsen:2019:LTQ

[493] Mikkel Slot Nielsen and Jan Pedersen. Limit theorems for quadratic forms and related quantities of discretely sampled continuous-time moving averages. ESAIM: Probability and Statistics, 23 (??):803–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Buraczewski:2019:LDE

[494] Dariusz Buraczewski and Mariusz Maślanka. Large deviation estimates for branching random walks. ESAIM:

Probability and Statistics, 23(??):823–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Godichon-Baggioni:2019:ASR

[495] Antoine Godichon-Baggioni. L^{p} and almost sure rates of convergence of averaged stochastic gradient algorithms: locally strongly convex objective. ESAIM: Probability and Statistics, 23(??):841–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Zheng:2019:PTT

[496] Guangqu Zheng. A Peccati–Tudor type theorem for Rademacher chaoses. ESAIM: Probability and Statistics, 23 (??):874–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chraibi:2019:OIP

[497] H. Chraibi, A. Dutfoy, T. Galtier, and J. Garnier. On the optimal importance process for piecewise deterministic Markov process. ESAIM: Probability and Statistics, 23(??):893–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Giraudo:2019:DIB

[498] Davide Giraudo. Deviation inequalities for Banach space valued martingales differences sequences and random fields. ESAIM: Probability and Statistics, 23 (??):922–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gugushvili:2019:BWN

[499] Shota Gugushvili, Frank van der Meulen, Moritz Schauer, and Peter Spreij. Bayesian wavelet de-noising with the caravan prior. ESAIM: Probability and Statistics, 23(??):947–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tanguy:2019:IOS

[500] Kevin Tanguy. Improved one-sided deviation inequalities under regularity assumptions for product measures. ESAIM: Probability and Statistics, 23 (??):979–??, ???? 2019. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dang:2020:EMF

[501] Thi-To-Nhu Dang. Estimation of the multifractional function and the stability index of linear multifractional stable processes. ESAIM: Probability and Statistics, 24(??):1–??, ???? 2020. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tang:2020:BEB

[502] Xufei Tang, Xuejun Wang, Yi Wu, and Fei Zhang. The Berry–Esseen bound of a wavelet estimator in non-randomly designed nonparametric regression model based on ANA errors. ESAIM: Probability and Statistics, 24(??):21–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Arbel:2020:SSG

[503] Julyan Arbel, Olivier Marchal, and Hien D. Nguyen. On strict sub-Gaussianity, optimal proxy variance and symmetry for bounded random variables. ESAIM: Probability and Statistics, 24(??):39–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Szablowski:2020:GKM

[504] Paweł J. Szabłowski. On the generalized Kesten–McKay distributions. ESAIM: Probability and Statistics, 24(??):56–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Puchkin:2020:AMN

[505] Nikita Puchkin and Vladimir Spokoiny.

An adaptive multiclass nearest neighbor classifier. ESAIM: Probability and Statistics, 24(??):69–??, ???? 2020. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mena:2020:CTM

[506] Ramsés H. Mena and Freddy Palma.

Continuous-time Markov processes, orthogonal polynomials and Lancaster probabilities. ESAIM: Probability and Statistics, 24(??):100–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Baldi:2020:TET

[507] Paolo Baldi. Tightness and exponential tightness of Gaussian probabilities. ESAIM: Probability and Statistics, 24 (??):113–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Guillotin-Plantard:2020:EPR

[508] Nadine Guillotin-Plantard, Françoise Pène, and Martin Wendler. Empirical processes for recurrent and transient random walks in random scenery. ESAIM: Probability and Statistics, 24 (??):127–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cerqueira:2020:NPS

[509] Andressa Cerqueira, Aurélien Garivier, and Florencia Leonardi. A note on perfect simulation for exponential random graph models. ESAIM: Probability and Statistics, 24(??):138–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pardoux:2020:LDE

[510] Etienne Pardoux and Brice Samegni-Kepgnou. Large deviations of the exit measure through a characteristic boundary for a Poisson driven SDE. ESAIM: Probability and Statistics, 24(??):148–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rackauskas:2020:BKI

[511] Alfredas Račkauskas and Charles Suquet. On Bernstein–Kantorovich invariance principle in Hölder spaces and weighted scan statistics. ESAIM: Probability and Statistics, 24(??):186–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Xiao:2020:PAQ

[512] Lishun Xiao, Shengjun Fan, and Dejian Tian. A probabilistic approach to quasilinear parabolic PDEs with obstacle and Neumann problems. ESAIM: Probability and Statistics, 24(??):207–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Clarenne:2020:RWD

[513] Adrien Clarenne. Rescaled weighted determinantal random balls. ESAIM: Probability and Statistics, 24(??):227–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Piccioni:2020:LSB

[514] Mauro Piccioni, Bartosz Kołodziejek, and Gérard Letac. Location and scale behaviour of the quantiles of a natural exponential family. ESAIM: Probability and Statistics, 24(??):244–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

DiBernardino:2020:EAP

[515] Elena Di Bernardino, Anne Estrade, and Maurizia Rossi. On the excursion area of perturbed Gaussian fields. ESAIM: Probability and Statistics, 24(??):252–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Doumas:2020:LZL

[516] Aristides V. Doumas and Vassilis G. Papanicolaou. The logarithmic Zipf law in a general urn problem. ESAIM: Probability and Statistics, 24(??):275–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Abraham:2020:VFG

[517] Romain Abraham, Aymen Bouaziz, and Jean-François Delmas. Very fat geometric Galton–Watson trees. ESAIM: Probability and Statistics, 24(??):294–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Olenko:2020:NCL

[518] Andriy Olenko and Volodymyr Vaskovych.

Non-central limit theorems for functionals of random fields on hypersurfaces. ESAIM: Probability and Statistics, 24 (??):315–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kwon:2020:ULC

[519] Younghak Kwon and Georg Menz. Uniform LSI for the canonical ensemble on the 1D-lattice with strong, finiterange interaction. ESAIM: Probability and Statistics, 24(??):341–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Boivin:2020:LDB

[520] Daniel Boivin and Thi Thu Hien Lê. Large deviations for Brownian motion in a random potential. ESAIM: Probability and Statistics, 24(??):374–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tugaut:2020:ETM

[521] Julian Tugaut. Exit-time of mean-field particles system. ESAIM: Probability and Statistics, 24(??):399–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kloeckner:2020:EMR

[522] Benoît R. Kloeckner. Empirical measures: regularity is a counter-curse to dimensionality. ESAIM: Probability and Statistics, 24(??):408–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Albertus:2020:RRE

[523] Mickael Albertus. Raking-ratio empirical process with auxiliary information learning. ESAIM: Probability and Statistics, 24(??):435–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Avram:2020:SFK

[524] Florin Avram, Danijel Grahovac, and Ceren Vardar-Acar. The W, Z scale functions kit for first passage problems of spectrally negative Lévy processes, and applications to control problems. ESAIM: Probability and Statistics, 24 (??):454–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fredes:2020:IMI

[525] Luis Fredes and Jean-François Marckert.

Invariant measures of interacting particle systems: Algebraic aspects. ESAIM: Probability and Statistics, 24(??):526–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mirakhmedov:2020:PLD

[526] Sherzod M. Mirakhmedov. The probabilities of large deviations for a certain class of statistics associated with multinomial distribution. ESAIM: Probability and Statistics, 24(??):581–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic). See correction [545].

Konzou:2020:ASE

[527] Essomanda Konzou and Angelo Efoevi Koudou. About the Stein equation for the generalized inverse Gaussian and Kummer distributions. ESAIM: Probability and Statistics, 24(??):607–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

DeBortoli:2020:RGR

[528] Valentin De Bortoli, Agnès Desolneux, Bruno Galerne, and Arthur Leclaire. Redundancy in Gaussian random fields.

ESAIM: Probability and Statistics, 24 (??):627–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ocafrain:2020:QSO

[529] William Ocafrain. Quasi-stationarity for one-dimensional renormalized Brownian motion. ESAIM: Probability and Statistics, 24(??):661–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Beyhum:2020:IRO

[530] Jad Beyhum. Inference robust to outliers with ℓ_{1}-norm penalization. ESAIM: Probability and Statistics, 24(??):688–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Alfonsi:2020:SQW

[531] Aurélien Alfonsi and Benjamin Jourdain. Squared quadratic Wasserstein distance: optimal couplings and Lions differentiability. ESAIM: Probability and Statistics, 24(??):703–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Vo:2020:CRS

[532] Thi Phuong Thuy Vo. Chain-referral sampling on stochastic block models. ESAIM: Probability and Statistics, 24 (??):718–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

O:2020:FBS

[533] Hun O, Mun-Chol Kim, and Chol-Kyu Pak. A framework of BSDEs with stochastic Lipschitz coefficients. ESAIM: Probability and Statistics, 24 (??):739–??, ???? 2020. CODEN ????

ISSN 1292-8100 (print), 1262-3318 (electronic).

Cholaquidis:2020:SEU

[534] Alejandro Cholaquidis and Antonio Cuevas. Set estimation under biconvexity restrictions. ESAIM: Probability and Statistics, 24(??):770–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Berger:2020:BMM

[535] Maxime Berger and Raphaël Cerf. A basic model of mutations. ESAIM: Probability and Statistics, 24(??):789–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Goehry:2020:RFT

[536] Benjamin Goehry. Random forests for time-dependent processes. ESAIM: Probability and Statistics, 24(??):801–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Brouste:2020:OSE

[537] Alexandre Brouste, Marius Soltane, and Irene Votsi. One-step estimation for the fractional Gaussian noise at highfrequency. ESAIM: Probability and Statistics, 24(??):827–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Azais:2020:SPE

[538] Jean-Marc Azaïs, François Bachoc, Agnès Lagnoux, and Thi Mong Ngoc Nguyen. Semi-parametric estimation of the variogram scale parameter of a Gaussian process with stationary increments. ESAIM: Probability and Statistics, 24(??):842–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gloter:2020:AID

[539] A. Gloter, I. Honoré, and D. Loukianova. Approximation of the invariant distribution for a class of ergodic jump diffusions. ESAIM: Probability and Statistics, 24(??):883–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Collins:2020:UCN

[540] Benoît Collins, Sushma Kumari, and Vladimir G. Pestov. Universal consistency of the k-NN rule in metric spaces and Nagata dimension. ESAIM: Probability and Statistics, 24(??):914–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Yao:2020:SRB

[541] Song Yao. L^{p} solutions of reflected backward stochastic differential equations with jumps. ESAIM: Probability and Statistics, 24(??):935–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Deschatre:2020:EFM

[542] Thomas Deschatre, Olivier Féron, and Marc Hoffmann. Estimating fast meanreverting jumps in electricity market models. ESAIM: Probability and Statistics, 24(??):963–??, ???? 2020. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Rommel:2021:QCS

[543] Cédric Rommel, J. Frédéric Bonnans, Baptiste Gregorutti, and Pierre Martinon. Quantifying the closeness to a set of random curves via the mean marginal likelihood. ESAIM: Probability and Statistics, 25(??):1–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bar-Lev:2021:NED

[544] Shaul K. Bar-Lev and Ad Ridder. New exponential dispersion models for count data: the ABM and LM classes. ESAIM: Probability and Statistics, 25(??):31–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mirakhmedov:2021:CPL

[545] Sherzod M. Mirakhmedov. Correction:

The probabilities of large deviations for a certain class of statistics associated with multinomial distributions. ESAIM: Probability and Statistics, 25(??):53–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic). See [526].

Nakarmi:2021:VBK

[546] Janet Nakarmi, Hailin Sang, and Lin Ge. Variable bandwidth kernel regression estimation. ESAIM: Probability and Statistics, 25(??):55–??, ???? 2021. CO-DEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Privault:2021:CER

[547] Nicolas Privault. Cardinality estimation for random stopping sets based on Poisson point processes. ESAIM: Probability and Statistics, 25(??):87–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dembin:2021:RTC

[548] Barbara Dembin. Regularity of the time constant for a supercritical Bernoulli percolation. ESAIM: Probability and Statistics, 25(??):109–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Popier:2021:BSV

[549] Alexandre Popier. Backward stochastic Volterra integral equations with jumps in a general filtration. ESAIM: Probability and Statistics, 25(??):133–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Coletti:2021:AST

[550] Cristian F. Coletti and Lucas R. de Lima. The asymptotic shape theorem for the frog model on finitely generated abelian groups. ESAIM: Probability and Statistics, 25(??):204–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Assaad:2021:WAS

[551] Obayda Assaad and Ciprian A. Tudor. Wavelet analysis for the solution to the wave equation with fractional noise in time and white noise in space. ESAIM: Probability and Statistics, 25(??):220–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cavalli:2021:PVL

[552] Benedetta Cavalli. A probabilistic view on the long-time behaviour of growthfragmentation semigroups with bounded fragmentation rates. ESAIM: Probability and Statistics, 25(??):258–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Basse-OConnor:2021:LAS

[553] Andreas Basse-O’Connor, Thorbjørn Grønbæk, and Mark Podolskij. Local asymptotic self-similarity for heavy tailed harmonizable fractional Lévy motions. ESAIM: Probability and Statistics, 25(??):286–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Obiang:2021:PGM

[554] Eunice Okome Obiang, Pascal Jézéquel, and Frédéric Proïa. A partial graphical model with a structural prior on the direct links between predictors and responses. ESAIM: Probability and Statistics, 25(??):298–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gaunt:2021:NEB

[555] Robert E. Gaunt. New error bounds for Laplace approximation via Stein’s method. ESAIM: Probability and Statistics, 25(??):325–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bansaye:2021:ERC

[556] Vincent Bansaye, Juan Carlos Pardo, and Charline Smadi. Extinction rate of continuous state branching processes in critical Lévy environments. ESAIM: Probability and Statistics, 25(??):346–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ceyhan:2021:LLN

[557] Elvan Ceyhan, John C. Wierman, and Pengfei Xiang. Law of large numbers for a two-dimensional class cover problem. ESAIM: Probability and Statistics, 25(??):376–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Debavelaere:2021:CEF

[558] Vianney Debavelaere and Stéphanie Allassonnière. On the curved exponential family in the Stochastic Approximation Expectation Maximization Algorithm. ESAIM: Probability and Statistics, 25(??):408–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cercle:2021:UBL

[559] Baptiste Cerclé. Unit boundary length quantum disk: a study of two different perspectives and their equivalence. ESAIM: Probability and Statistics, 25 (??):433–??, ???? 2021. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Journel:2022:CPA

[560] Lucas Journel and Pierre Monmarché. Convergence of a particle approximation for the quasi-stationary distribution of a diffusion process: Uniform estimates in a compact soft case. ESAIM: Probability and Statistics, 26(??):1–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dai:2022:ETA

[561] Hongshuai Dai, Donald A. Dawson, and Yiqiang Q. Zhao. Exact tail asymptotics for a three-dimensional Browniandriven tandem queue with intermediate inputs. ESAIM: Probability and Statistics, 26(??):26–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Budhiraja:2022:AQS

[562] Amarjit Budhiraja, Nicolas Fraiman, and Adam Waterbury. Approximating quasi-stationary distributions with interacting reinforced random walks. ESAIM: Probability and Statistics, 26 (??):69–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Amorino:2022:OCR

[563] Chiara Amorino and Eulalia Nualart.

Optimal convergence rates for the invariant density estimation of jump-diffusion processes. ESAIM: Probability and Statistics, 26(??):126–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Logachov:2022:LTC

[564] A. Logachov, A. Mogulskii, and A. Yambartsev. Limit theorems for chains with unbounded variable length memory which satisfy Cramer condition. ESAIM: Probability and Statistics, 26(??):152–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Vu:2022:RCG

[565] Mai Trang Vu, François Bachoc, and Edouard Pauwels. Rate of convergence for geometric inference based on the empirical Christoffel function. ESAIM: Probability and Statistics, 26(??):171–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mantoux:2022:AAM

[566] Clément Mantoux, Stanley Durrleman, and Stéphanie Allassonnière. Asymptotic analysis of a matrix latent decomposition model. ESAIM: Probability and Statistics, 26(??):208–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Sun:2022:MSS

[567] Chengfeng Sun, Hongjun Gao, Hui Liu, and Jie Zhang. Martingale solutions of the stochastic 2D primitive equations with anisotropic viscosity. ESAIM: Probability and Statistics, 26(??):243–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Ekstrom:2022:BSC

[568] Erik Ekström and Yuqiong Wang. Bayesian sequential composite hypothesis testing in discrete time. ESAIM: Probability and Statistics, 26(??):265–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Joly:2022:TNA

[569] Émilien Joly and Bastien Mallein. A tractable non-adaptative group testing method for non-binary measurements. ESAIM: Probability and Statistics, 26 (??):283–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Chen:2022:PRI

[570] Junchao Chen, Noufel Frikha, and Houzhi Li. Probabilistic representation of integration by parts formulae for some stochastic volatility models with unbounded drift. ESAIM: Probability and Statistics, 26(??):304–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bogdanskii:2022:LBR

[571] Victor Bogdanskii, Ilya Pavlyukevich, and Andrey Pilipenko. Limit behaviour of random walks on Z m with twosided membrane. ESAIM: Probability and Statistics, 26(??):352–??, ???? 2022.

CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bednorz:2022:OUA

[572] Witold M. Bednorz, Rafał M. Łochowski, and Rafał Martynek. On optimal uniform approximation of Lévy processes on Banach spaces with finite variation processes. ESAIM: Probability and Statistics, 26(??):378–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Tomasevic:2022:EBM

[573] Milica Tomašević, Vincent Bansaye, and Amandine Véber. Ergodic behaviour of a multi-type growth-fragmentation process modelling the mycelial network of a filamentous fungus. ESAIM: Probability and Statistics, 26(??):397–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Backhoff-Veraguas:2022:BLW

[574] Julio Backhoff-Veraguas, Joaquin Fontbona, Gonzalo Rios, and Felipe Tobar. Bayesian learning with Wasserstein barycenters. ESAIM: Probability and Statistics, 26(??):436–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Mirakhmedov:2022:IAE

[575] Sherzod M. Mirakhmedov. On the intermediate asymptotic efficiency of goodness-of-fit tests in multinomial distributions. ESAIM: Probability and Statistics, 26(??):473–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Jourdain:2022:ODM

[576] B. Jourdain and W. Margheriti. One dimensional martingale rearrangement couplings. ESAIM: Probability and Statistics, 26(??):495–??, ???? 2022. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gradinaru:2023:ABT

[577] Mihai Gradinaru and Emeline Luirard. Asymptotic behavior for a timeinhomogeneous Kolmogorov type diffusion. ESAIM: Probability and Statistics, 27(??):1–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hamaguchi:2023:AAM

[578] Yushi Hamaguchi and Dai Taguchi. Approximations for adapted m-solutions of type-II backward stochastic Volterra integral equations. ESAIM: Probability and Statistics, 27(??):19–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Duchemin:2023:RPM

[579] Quentin Duchemin. Reliable prediction in the Markov stochastic block model. ESAIM: Probability and Statistics, 27 (??):80–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Liu:2023:SRP

[580] Chong Liu, David J. Prömel, and Josef Teichmann. A Sobolev rough path extension theorem via regularity structures. ESAIM: Probability and Statistics, 27(??):136–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Hariz:2023:FCW

[581] Samir Ben Hariz, Alexandre Brouste, Youssef Esstafa, and Marius Soltane. Fast calibration of weak Farima models. ESAIM: Probability and Statistics, 27(??):156–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Blath:2023:MVE

[582] Jochen Blath and András Tóbiás. Microbial virus epidemics in the presence of contact-mediated host dormancy. ESAIM: Probability and Statistics, 27(??):174–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Louvet:2023:SMV

[583] Apolline Louvet. Stochastic measurevalued models for populations expanding in a continuum. ESAIM: Probability and Statistics, 27(??):221–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Monmarche:2023:ECA

[584] Pierre Monmarché. Elementary coupling approach for non-linear perturbation of Markov processes with mean-field jump mechanisms and related problems. ESAIM: Probability and Statistics, 27 (??):278–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Bourquin:2023:PRS

[585] Antoine Bourquin. Persistence in randomly switched Lotka–Volterra food chains. ESAIM: Probability and Statistics, 27(??):324–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Pang:2023:MPE

[586] Guodong Pang and Étienne Pardoux. Multi-patch epidemic models with general exposed and infectious periods. ESAIM: Probability and Statistics, 27 (??):345–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Lecestre:2023:REF

[587] Alexandre Lecestre. Robust estimation in finite mixture models. ESAIM: Probability and Statistics, 27(??):402–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dumbgen:2023:SOT

[588] Lutz Dümbgen and Alexandre Mösching. On stochastic orders and total positivity. ESAIM: Probability and Statistics, 27(??):461–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Godichon-Baggioni:2023:NAA

[589] Antoine Godichon-Baggioni, Nicklas Werge, and Olivier Wintenberger. Nonasymptotic analysis of stochastic approximation algorithms for streaming data. ESAIM: Probability and Statistics, 27(??):482–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Miclo:2023:SST

[590] Laurent Miclo. Strong stationary times for finite Heisenberg walks. ESAIM: Probability and Statistics, 27(??):515–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Xu:2023:MDP

[591] Jie Xu, Jiayin Gong, and Jie Ren.

A moderate deviation principle for stochastic Hamiltonian systems. ESAIM: Probability and Statistics, 27(??):558–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Minh:2023:SBE

[592] Tâm Le Minh. U-statistics on bipartite exchangeable networks. ESAIM: Probability and Statistics, 27(??):576–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Varet:2023:NPP

[593] Suzanne Varet, Claire Lacour, Pascal Massart, and Vincent Rivoirard. Numerical performance of penalized comparison to overfitting for multivariate kernel density estimation. ESAIM: Probability and Statistics, 27(??):621–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Etore:2023:PPV

[594] P. Etoré, J. R. León, and C. Prieur.

A probabilistic point of view for the Kolmogorov hypoelliptic equations. ESAIM: Probability and Statistics, 27 (??):668–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Massing:2023:AEA

[595] Till Massing. Approximation and error analysis of forward–backward SDEs driven by general Lévy processes using shot noise series representations. ESAIM: Probability and Statistics, 27 (??):694–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

deRaynal:2023:RET

[596] Paul-Eric Chaudru de Raynal, Manh Hong Duong, Pierre Monmarché, Milica Tomašević, and Julian Tugaut. Reducing exit-times of diffusions with repulsive interactions. ESAIM: Probability and Statistics, 27(??):723–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Fournier:2023:CEM

[597] Nicolas Fournier. Convergence of the empirical measure in expected Wasserstein distance: non-asymptotic explicit bounds in R^{d}. ESAIM: Probability and Statistics, 27(??):749–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Durand:2023:WSS

[598] Amaury Durand and François Roueff. Weakly stationary stochastic processes valued in a separable Hilbert space: Gramian–Cramér representations and applications. ESAIM: Probability and Statistics, 27(??):776–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Albin:2023:ABS

[599] J. M. P. Albin and Mattias Sundén. On the asymptotic behaviour of superexponential Lévy processes. ESAIM: Probability and Statistics, 27(??):810–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Brehier:2023:AID

[600] Charles-Edouard Brehier. Approximation of the invariant distribution for a class of ergodic SDEs with one-sided Lipschitz continuous drift coefficient using an explicit tamed Euler scheme. ESAIM: Probability and Statistics, 27 (??):841–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Velleret:2023:EQE

[601] Aurélien Velleret. Exponential quasiergodicity for processes with discontinuous trajectories. ESAIM: Probability and Statistics, 27(??):867–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Peng:2023:NFG

[602] Cheng Peng. A new formulation of generalized gamma: Some results and applications. ESAIM: Probability and Statistics, 27(??):913–??, ???? 2023. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

O:2024:WSO

[603] Hun O, Mun-Chol Kim, and Kon-Gun Kim. Wellposedness of second order reflected BSDEs: A new formulation. ESAIM: Probability and Statistics, 28 (??):1–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Seguret:2024:DAM

[604] Adrien Séguret, Thomas Le Corre, and Nadia Oudjane. A decentralized algorithm for a mean field control problem of piecewise deterministic Markov processes. ESAIM: Probability and Statistics, 28(??):22–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Aleksian:2024:SID

[605] A. Aleksian, P. Del Moral, A. Kurtzmann, and J. Tugaut. Self-interacting diffusions: Long-time behaviour and exit-problem in the uniformly convex case. ESAIM: Probability and Statistics, 28(??):46–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Gapeev:2024:CCP

[606] Pavel V. Gapeev and Monique Jeanblanc. On the construction of conditional probability densities in the Brownian and compound Poisson filtrations. ESAIM: Probability and Statistics, 28 (??):62–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Vovchanskyi:2024:SSC

[607] M. B. Vovchanskyi. Splitting for some classes of homeomorphic and coalescing stochastic flows. ESAIM: Probability and Statistics, 28(??):75–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Protter:2024:STO

[608] Philip Protter and Alejandra Quintos. Stopping times occurring simultaneously. ESAIM: Probability and Statistics, 28(??):110–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Kumari:2024:UCK

[609] Sushma Kumari and Vladimir G. Pestov. Universal consistency of the k-NN rule in metric spaces and Nagata dimension. II. ESAIM: Probability and Statistics, 28(??):132–??, ???? 2024.

CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Cavallazzi:2024:IKF

[610] Thomas Cavallazzi. Itô–Krylov’s formula for a flow of measures. ESAIM: Probability and Statistics, 28(??):161–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Langholm:2024:DND

[611] Tore Langholm and Harald Totland. Deformed normal distributions. ESAIM: Probability and Statistics, 28(??):195–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).

Dombry:2024:VLR

[612] Clément Dombry and Youssef Esstafa.

The vanishing learning rate asymptotic for linear L^{2}-boosting. ESAIM: Probability and Statistics, 28(??):227–??, ???? 2024. CODEN ???? ISSN 1292-8100 (print), 1262-3318 (electronic).